### v1

# Path-integral Monte Carlo estimator for the dipole polarizability of quantum plasma

Abstract
We present a path-integral Monte Carlo estimator for calculating the dipole polarizability of interacting Coulomb plasma in the long-wavelength limit, i.e., the optical region. Unlike the conventional dynamic structure factor in reciprocal space, our approach is based on the real-space dipole autocorrelation function and is suited for long wavelengths and small cell sizes, including finite clusters. The simulation of thermal equilibrium in imaginary time has exact Coulomb interactions and Boltzmann quantum statistics. For reference, we demonstrate analytic continuation of the Drude model into the imaginary time and Matsubara series, showing perfect agreement with our data within ranges of finite temperatures and densities. Method parameters, such as the finite time-step and finite-size effects prove only modestly significant. Our method, here carefully validated against an exactly solvable reference, remains amenable to more interesting domains in higher-order optical response, quantum confinements and quantum statistical effects, and applications in plasmonics, heterogeneous plasmas and nonlinear optics, such as epsilon-near-zero materials.
preprint: APS/123-QED
I Introduction
The Drude model, introduced at the beginning of the last century by Paul Drude as an attempt at a comprehensive electron theory of metals [1, 2], remains nowadays ubiquitous in describing the optical response of metals, as well as optical materials with metal-like behaviours, such as epsilon-near-zero (ENZ) media [3]. The Drude parameters have been fitted to experiments [4] and hydrodynamic models of various kinds typically recover the Drude model or something close to it [5]. One of the advantages of the Drude model, beside its simplicity, resides in its high effectiveness in describing the optical response in a broad range of frequencies, from the near IR to the UV regime. This allows to characterize the behaviour of different photonic materials, such as metals [6, 7], plasmonic circuits [8], ENZ materials [9], as well as more complicated photonics platforms, such as metasurfaces [10], with an effective Drude model.
Thus, there is great intrigue towards the recovery of the effective response properties from first principles. In spirit, the Drude model describes the dynamics of non-interacting free electrons, which occasionally scatter from ions. When the scattering is rare, this model becomes the homogeneous electron gas (HEG), or jellium, whose quantum mechanical linear response function is the Lindhard function $\chi_{0}(\mathbf{q},\omega))$ [11], where $\mathbf{q}$ and $\omega$ treat, respectively, the momentum and the frequency of the incoming field. Not surprisingly, the Lindhard function, despite begin famously hard to evaluate generally, recovers the Drude model upon taking approximations of long wavelength, high temperature and low scattering [7, 6].
The Lindhard response can be expressed more generally as [12, 13]
$\chi(\mathbf{q},\omega)=\frac{\chi_{0}(\mathbf{q},\omega)}{1+v_{q}(1-G(\mathbf{q},\omega))\chi_{0}(\mathbf{q},\omega)},$ (1)
where the fraction treats dielectric screening, $v_{q}=e^{2}/q^{2}\varepsilon_{0}$ is the reciprocal Coulomb potential (in atomic units) and $G(\mathbf{q},w)$ is the so-called local field correction, basically treating effects due to Fermion exchange and dynamic correlations. For decades, numerous works (e.g. Refs. [12, 14, 13]) have discussed and featured either static $G(\mathbf{q},\omega=0)$ or dynamic $G(\mathbf{q},\omega)$ in pursuit of the non-trivial response within the so-called regime of warm, dense matter (WDM).
Recently, significant progress has been made to calculate the dielectric response of the WDM with the path-integral Monte Carlo method (PIMC) [15, 16, 17, 18, 19, 13, 20, 21]. The PIMC approach can treat accurate quantum many-body interactions at non-trivial temperatures, and it is in principle only subject to numerical challenges due to the Fermion sign problem (FSP) [22, 23]. Within this framework, moreover, the dielectric response is directly connected to the dynamic structure factor through the intermediate scattering function [11]. The latter can be straightforwardly estimated from the PIMC walker as the density-density correlation function in the imaginary time. This function can be then analytically continued in the complex plane to recover the full dynamic response of the system. The inverse transformation back to the real time, however, faces serious challenges with numerical noise, prevalent in Monte Carlo simulation, and thus requires special methods [24, 25, 20].
Within this framework, Dornheim et al. have comprehensively studied the first [13] and higher order [17, 18] response properties of the HEG. An apparent limitation of their approach is that the allowed values for the wavenumber are limited by $q=2\pi n/L$, where $n\geq 1$ and $L$ is the (cubic) simulation cell size. Therefore, the typical optical wavelengths with $q\ll L^{-1}$, thus requiring large $L$, are hard to accomplish because of the algorithmic scaling and performance loss attributed to the high particle count $N$. Indeed, the literature focuses on characterizing inelastic neutron scattering or X-ray Thompson scattering spectra, where the momentum transfers are significant [13] compared to that of photons.
Here, we explore a complementary PIMC approach to regain focus on the long, optical wavelengths. Namely, we demonstrate computation of the conventional optical response based on the complex dipole polarizability, or the dipole-dipole correlation function. This approach, which evades the difficulty of low $q$, has been studied before in finite systems [26, 27, 28, 29] but not in periodic ones. Granted, formal problems arise that are more naturally addressed in the Fourier space: The long-wavelength limit is only quasi-accurate, since we treat optical fields that are locally constant inside a finite box, i.e. $\lambda\gg L$, and formally expand that by summing over infinite periodic images. In practice, there is no field in the simulation, and thus, the approximation remains justified in all parameter ranges in the intended scope: for example, the box size $L$ required for $N=16$ particles to simulate the metallic density of gold is over $300$ smaller than an ultraviolet wavelength $\lambda=200$ nm.
However, it is not obvious, a priori, how a PIMC estimator for the dipole polarizability must be laid out for periodic systems or how it responds to finite-size effects or converge parameters, such as the finite time-step. Therefore, it is the aim of this work to establish these matters in a controllable setting: We first discuss analytical correspondence between the Drude model and the PIMC polarizability, evaluated in the imaginary time. For convenience and consistency with the Drude model, we use distinguishable particles, i.e. the Boltzmann statistics, to evade the Fermion sign problem—for now. We then survey various methodological and physical parameters to assert that the correspondence holds, when the Drude scattering rate is low. Indeed, we show that biases are statistically non-distinguishable within $2\%$ accuracy; the PIMC polarizability matches the Drude model. Formally, this is well anticipated, but practically the lack of biases is a signal of a robust simulation method. The assessments of performance help us avoid pitfalls, establish good practices and assert validity for continuation studies.
This work is organised as follows: In Sec. II, we lay out the theoretical context of the estimator and the PIMC simulation. In Sec. III we overview the computational details. In Sec. IV we discuss the numerical results and in Sec. V we summarize the work and future prospects of this methodology.
II Theory
In this section, we first overview the linear optical response of quantum particles in the long-wavelength limit. We then consider analytic continuation of the Drude model into imaginary domain, where it meets the dipole moment autocorrelation function accessible to the PIMC method. Throughout this work, we use the Hartree atomic units: $\hbar=1/(4\pi\epsilon_{0})=m_{e}=1$.
II.1 Electric field response
The first-order local polarization $\mathbf{P}(\mathbf{r},\omega)$ of charged plasma responding to an incident electric field $\mathbf{F}(\mathbf{r^{\prime}},\omega)$ can be written as
$\mathbf{P}(\mathbf{r},\omega)=\varepsilon_{0}\chi(\mathbf{r},\mathbf{r^{\prime}},\omega)\mathbf{F}(\mathbf{r^{\prime}},\omega),$ (2)
where $\chi(\mathbf{r},\mathbf{r^{\prime}},\omega)$ is a complex linear response function, the electric susceptibility, in the local response approximation. Let us consider the long-wavelength limit, in which the field is oscillating with a wavelength $\lambda=2\pi c/\omega\gg L$, where $c$ is the speed of light and $L$ is a measure of longitudinal distance containing the interaction. Then, within $L$ we may consider a locally constant field
$\mathbf{F}(\mathbf{r^{\prime}},\omega)\rightarrow\delta(\mathbf{r^{\prime}}-\mathbf{r})\mathbf{F}(\omega),$ (3)
where $\delta$ is the Dirac delta function, inducing total mean polarization $P(\omega)\equiv\varepsilon_{0}\chi(\omega)\mathbf{F}(\omega)$ according to total mean susceptibility
$\chi(\omega)=\frac{1}{L^{d}}\int_{0}^{L}\chi(\mathbf{r},\mathbf{r^{\prime}},\omega)\delta(\mathbf{r}-\mathbf{r^{\prime}})d\mathbf{r^{\prime}},$ (4)
where $d$ is the dimensionality of $\mathbf{r}$. We may further consider the per-particle dipole polarizability $\alpha(\omega)$ through
$\varepsilon_{0}\chi(\omega)=n\alpha(\omega),$ (5)
where $n=N/L^{3}$ is the number density of the particles within $L$. The tensorial characters of $\chi$ and $\alpha$ are omitted here for simplicity of notation; because of the radial symmetry of the plasma, we only consider the diagonal, where $\mathbf{P}$ and $\mathbf{F}$ are parallel.
For quantized plasma represented by $N$ particles, the per-particle dipole polarizability $\alpha(\omega)$ can also be expressed using the Kubo formula [30]
$\alpha(\omega)$ $=-\frac{i}{N\hbar}\int_{0}^{\infty}\mathrm{d}te^{i\omega t}\int d\mathbf{r}\langle[\mu(\mathbf{r},t),\mu(\mathbf{r},0)]\rangle$ (6)
$\equiv-\int_{-\infty}^{\infty}\mathrm{d}te^{i\omega t}G^{R}_{\mu\mu}(t)$ (7)
$\equiv-G^{R}_{\mu\mu}(\omega),$ (8)
where $G^{R}$ denotes the retarded dipole-dipole correlation function and the dipole moment operator is given by
$\mu(\mathbf{r},t)=\sum_{i=1}^{N}q_{i}[\delta(\mathbf{r}-\mathbf{r_{i}}(t))-\langle\mathbf{r_{i}}\rangle],$ (9)
where $q_{i}$ is the charge and $\mathbf{r_{i}}(t)$ the time-dependent displacement of the $i$th particle, and $\langle\mathbf{r_{i}}\rangle$ is their mean displacement, a non-perturbed reference point. Clearly, when $\sum q_{i}\neq 0$ the value depends on the reference point, which is problematic in translationally invariant systems, such as plasmas. Therefore, it is more meaningful to consider the relative correlation function
$\tilde{G}^{R}_{\mu\mu}(t)\equiv G^{R}_{\mu\mu}(t)-G^{R}_{\mu\mu}(0)$ (10)
which is zero at $t=0$ but nevertheless yields
$\tilde{\alpha}(\omega)$ $\equiv-\int_{-\infty}^{\infty}\mathrm{d}te^{i\omega t}\tilde{G}^{R}_{\mu\mu}(t)=-G^{R}_{\mu\mu}(\omega)=\alpha(\omega)$ (11)
for all $\omega\neq 0$, while $\alpha(\omega=0)$ remains elusive. On the other hand, it is well known that the total response of charged gas in the limits of $\omega=0$ and $q=0$ (long wavelength) is zero due to perfect screening [31]. The linear response framework considered here does not treat screening, but this should justify why $\alpha(\omega=0)$ is unimportant for the free gas.
The relative polarizability $\tilde{\alpha}(\omega)$ is analytic in the upper complex plane and can be analytically continued to imaginary time $it\rightarrow\tau$. The associated, time-ordered imaginary-time correlation function is given by
$G_{\mu\mu}(\tau)=\langle\mathcal{T}\mu(0)\mu(\tau)\rangle,$ (12)
where $\mathcal{T}$ is the time-ordering operator. The analytic continuation can be done via the spectral density function
$A(\omega)\equiv-2\mathrm{Im}[G^{R}_{\mu\mu}(\omega)]$ (13)
where the bosonic integral transformation yields
$G_{\mu\mu}(\tau)=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\frac{\mathrm{e}^{-\tau\omega}}{1-\mathrm{e}^{-\beta\omega}}A(\omega),$ (14)
where $G_{\mu\mu}(\tau)=G_{\mu\mu}(\tau+n\beta)$ for any integer $n$ and $\beta=1/k_{B}T$ is the inverse temperature. It is easy to see that
$\tilde{G}_{\mu\mu}(\tau)$ $\equiv G_{\mu\mu}^{R}(\tau)-G_{\mu\mu}^{R}(0)$ (15)
$=-\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\frac{1-\mathrm{e}^{-\tau\omega}}{1-\mathrm{e}^{-\beta\omega}}A(\omega).$ (16)
Alternatively, we may write
$G_{\mu\mu}(i\omega_{n})=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\frac{1}{i\omega_{n}-\omega}A(\omega),$ (17)
where $\omega_{n}=2\pi n/\hbar\beta$ are the bosonic Matsubara frequencies.
While the imaginary-domain functions $G_{\mu\mu}(\tau)$ and $G_{\mu\mu}(i\omega_{n})$ contain the same information as $G^{R}_{\mu\mu}(\omega)$, the conversion from imaginary to real domain by inverting Eqs. (14) or (17) is a known ill-posed problem, which is accentuated by the unavoidable error bars present in Quantum Monte Carlo methods [24]. The use of existing numerical methods of analytic continuation, such as Maximum Entropy (MaxEnt) [25, 26], will be left for future consideration.
II.2 Analytic continuation of the Drude model
The ubiquitous model for considering the optical response of microscopic particles is the Drude model. The Drude susceptibility is given by
$\chi^{D}(\omega)=-\frac{\omega_{p}^{2}}{\omega^{2}+i\omega\Gamma},$ (18)
where $\omega_{p}^{2}=\tfrac{ne^{2}}{m\varepsilon_{0}}=\tfrac{e^{2}}{4\pi\varepsilon_{0}m}\tfrac{3}{r_{s}^{3}}$ is the plasma frequency and $\Gamma$ is a positive damping coefficient. The Drude per-particle polarizability is
$\alpha^{D}(\omega)=\frac{\chi_{e}^{D}(\omega)\varepsilon_{0}}{n},$ (19)
whose spectral density, according to Eq (13), is given by
$A^{D}(\omega)$ $=\frac{2\omega_{p}^{2}\varepsilon_{0}}{n}\frac{1}{\Gamma\omega(1+\omega^{2}/\Gamma^{2})}$ (20)
$=\frac{2e^{2}}{m}\frac{1}{\Gamma\omega(1+\omega^{2}/\Gamma^{2})}.$ (21)
The spectrum diverges at $\omega=0$ and its analytic continuation with Eq. (14) becomes problematic. However, the relative polarizability $\tilde{\alpha}(\omega)$ can be obtained from Eq. (16) because the divergence is controlled at $\omega=0$ by the transformation kernel.
Furthermore, an analytical expression for $G^{D}_{\mu\mu}(i\omega_{n})$, which is the correlation function in Matsubara frequencies associated to the Drude model, can be obtained by using Eq.  (20) and substituting into Eq. 17. This yields
$G^{D}_{\mu\mu}(i\omega_{n})=\frac{2\pi e^{2}}{m}\left[\frac{1}{\omega_{n}^{2}/\Gamma+\omega_{n}}\right].$ (22)
The analytical expression given by Eq.  (22) is valid for $\omega_{n}>0$. The static limit is defined here based on the zero-frequency Fourier transform of relative correlation function
$\tilde{G}^{D}_{\mu\mu}(i\omega_{n})=\int_{0}^{\beta}d\tau\tilde{G}^{D}_{\mu\mu}(\tau),$ (23)
but as discussed before, it is not physical and only used for numerical assessment of performance.
From Eq. (20) it is clear that the Drude spectrum does not depend on the physical density, because the oscillators are not interacting; rather, the density only shows in the transformation to the macroscopic susceptibility with Eq.(5). To consider the interacting problem, let us therefore turn to numerical path-integral simulation.
II.3 PIMC estimator
The PIMC method can be used to simulate interacting quantum particles at finite temperatures. The method draws samples from the distribution of the thermal density matrix:
$\rho(R,R;\beta)=Z^{-1}\int dR\langle R|\hat{\rho}(\beta)|R\rangle,$ (24)
where $R$ denotes a set of variable particle coordinates, $\hat{\rho}(\beta)=e^{-\beta H}$ is the density operator at inverse temperature $\beta$, $H$ is the Hamiltonian and $Z=\langle R|\hat{\rho}|R\rangle$ is the thermal partition function. Equation (24) overlooks summations arising from spin degrees of freedom and particle permutations required to treat proper quantum statistics of bosons or fermions; instead, the simulation treats so-called boltzmannons, which comprise a Coulomb interacting gas of distinguishable quantum particles. The boltzmannons evade the FSP but do not correspond to physical electrons, alas, except at high temperatures and low densities [23] (or strong coupling [32]), where particle repulsion outweighs exchange and the quantum statistics merge. Our results are also affected by the (lack of proper) quantum statistics.
The interacting Hamiltonian for $N$ distinguishable particles is given by
$H(R)=T+V=\lambda_{m}\sum_{i}\nabla_{i}^{2}+\sum_{i<j}\frac{q_{i}q_{j}}{|\mathbf{r}_{i}-\mathbf{r}_{j}|}$ (25)
where $\lambda_{m}=1/2m_{e}$ and $T$ and $V$ correspond, respectively, to the summation over kinetic and potential operators, the last ones comprising exact Coulomb interaction between pairs of particles. Since $T$ and $V$ do not commute, the density operator must be expanded with the semi-group property
$e^{-\beta(T+V)}=\lim_{M\rightarrow\infty}\left[e^{-\Delta\tau T}e^{-\Delta\tau V}\right],$ (26)
where $\Delta\tau=\beta/M$ is the time-step and $M$ is the Trotter number. Using Eq. (26) and the product property [23], the exact Coulomb action between pairs of species can be treated for finite time-steps [33]. In this so-called pair approximation, higher-body interactions retain a modest time-step error proportional to $(\Delta\tau)^{3}$ [34, 33]. Therefore, it is justified to choose finite $M$, making the PIMC integrand a discrete walker of $M$ times $R$ coordinates, which forms a closed loop, an imaginary-time trajectory:
$\rho(R,R;\beta)=Z^{-1}\prod_{j=1}^{M}dR_{j}\langle R_{j-1}|\hat{\rho}(\Delta\tau)|R_{j}\rangle,$ (27)
where $R_{j}$ contains the real-space coordinates for $N$ particles at $j$th time-slice, and the loop closes with $R_{M}=R_{0}$. The trajectory is efficiently sampled as a Markov chain using the Metropolis multilevel bisection algorithm, as described in [35, 23, 33].
The calculation of the thermal expectation values of diagonal properties, such as potential energy $V$, is done from Eq. (27) as
$\langle V\rangle$ $=Z^{-1}\int dR\langle R|\hat{V}\hat{\rho}(\beta)|R\rangle$ (28)
$=Z^{-1}\prod_{j=1}^{M}dR_{j}\rho(R_{j-1},R_{j};\Delta\tau)V(R_{0})$ (29)
$=\frac{1}{M}\sum_{j=1}^{M}\langle V(R_{j}),$ (30)
where the last line employs symmetry properties of the trace, enabling to calculate mean over all time-slices of any trajectory. Because of the kinetic operator, the total energy $E$ is more complicated; here we use the thermal estimator [23], whose numerically exact pair energy $\partial\rho/\partial\beta$ can be precomputed [33].
The calculation of time-ordered observables, such as the imaginary-time correlation function from Eq. (12), is equally straightforward. The mean per-particle correlation function is evaluated as
$\langle G_{\mu\mu}(\tau)\rangle=\frac{1}{ZN}\prod_{j=1}^{M}dR_{j}\rho(R_{j-1},R_{j};\Delta\tau)\mu(R_{0})\mu(R_{t}),$ (31)
where the time-difference index is $t=j+\tau/\Delta\tau$, modulo $M$. The expression can be easily symmetrized, optimized for evaluation and generalized for different combinations of measurement operators, as detailed in [28, 29]. In the path-integral context, the dipole moment operator is
$\mu(R_{t})$ $=\sum_{i=1}^{N}q_{i}\left[\mu(\mathbf{r}_{i}(t\Delta\tau)-\langle\mu_{i}\rangle\right],$ (32)
which again depends on the point of reference $\langle\mu_{i}\rangle$. However, as was done in Ref. [26], we identify
$\langle\tilde{G}_{\mu\mu}(\tau)\rangle$ $=\frac{1}{2N}\langle(\mu(0)-\mu(\tau))^{2}\rangle,$ (33)
where the relative correlation function arises from mean squared fluctuations of the instantaneous dipole moment along the imaginary-time path.
II.4 Periodic boundary conditions
The simulation of charged plasma requires periodic boundary conditions (PBC). Here we use a cubic box of size $L$, whose size is determined by
$L=r_{s}\left(\frac{4\pi N}{3}\right)^{(1/3)},$ (34)
where $r_{s}=(3/4\pi n)^{1/3}$ is the Wigner-Seitz radius of the particle gas. The treatment of periodic Coulomb images is treated using the ordinary Ewald summation [36, 37]. Note that the long-wavelength approximation discussed in Sec. II.1 also uses $L$, and the symbols are in practice interchangeable but semantically different.
The finite density is enforced for the many-body interactions of the plasma, but otherwise the evaluation resembles an open-boundary problem. So, the evaluation of the dipole polarizability estimator, Eq. (33), is done on walkers whose all particle paths are unfolded into continuous trajectories near boundaries. Since the simulated particles may not exchange paths (permute), non-zero winding of a path through the cell is an artifact of the finite size. Therefore, we reject all moves leading to non-zero winding. Beside this constraint, the sampling is done using the multilevel bisection method [23]. The probability of winding increases when the thermal wavelength of the bisection move approaches the cell dimension, i.e.
$\sqrt{2^{l}\lambda_{m}\Delta\tau}\sim L,$ (35)
where $l$ is the bisection multilevel. We constrain $l$ so that the l.h.s. of Eq. (35) is lower than $L/8$. This is a simplification to avoid using periodic free-particle density matrix [23, 33], which would not inhibit the winding, anyway. As a result, fewer than one winding move per $10^{3}$ trial moves is observed and has to be rejected in any density. A modest sampling bias is mitigated along other finite-size effects when $L$ naturally grows with $N^{1/3}$.
III Computational details
The PIMC simulation software, including the exact Coulomb pair action generation, implements the canonical Metropolis Monte Carlo algorithm for a PIMC walker in Fortran90, featuring parallel sampling and data binning of a number of Markovian walkers. Here, 40 CPU cores are used for parallel sampling. Raw data after sample binning is stored in hierarchical data format (HDF5) and data processing chains operate on standard numerical Python libraries. The integral transformations, such as Eq. (16), are carried out using numerical quadrature. The simulation campaigns are operated with Nexus workflows [38] on various high-performance computing (HPC) facilities listed in Sec. V. All implementation details, including code snapshots and workflow scripts for step-by-step reproduction of the simulations, postprocessing and individual figures are supplied in a separate data repository [39].
[…]
IV Results
Figure 1: Comparison overview of the relative dipole correlation function $\tilde{G}_{\mu\mu}$ from PIMC and the Drude model. In the top, the imaginary-time correlation function $\tilde{G}_{\mu\mu}(\tau)$ is plotted at two temperatures, $\Theta_{1}=0.1$ and $\Theta_{2}=0.2$, resulting in two curves ending at different values for $\beta_{1}=2\beta_{2}$. In the bottom, the first few corresponding Matsubara data $\tilde{G}_{\mu\mu}(i\omega_{n})$ are plotted, where the frequencies $\omega_{n}=2\pi n/\hbar\beta$ for $\Theta_{1}$ are twice those of $\Theta_{2}$, but the two functions lie on the same curve. In both cases, the data match the analytically continued Drude model within statistical uncertainties.
Let us now discuss results of the PIMC simulation, laid out in the earlier sections, as applied to charged quantum plasma at various finite temperatures and densities. With $m=m_{e}$ and $q^{2}=e^{2}$, the simulated particles resemble free electrons, only without sign-flipping permutations. Thus, we use here conventional units of the HEG and express densities with $r_{s}$ and finite temperatures with $\Theta=T/T_{f}$, where $k_{B}T_{f}=\hbar^{2}k_{f}^{2}/2$ and $k_{f}=(3n\pi^{2})^{1/3}$ are, respectively, the Fermi temperature and the Fermi wavevector. The covered densities range between $r_{s}=2.0\ldots 8$ a0 and finite temperatures between $\Theta=T/T_{f}=0.1\ldots 0.8$.
The main observables of interest are overviewed in Fig. 1: the relative imaginary-time correlation function $\tilde{G}_{\mu\mu}(\tau)$ and its Matsubara series $G_{\mu\mu}(i\omega_{n})$. As stated in Sec. II, the vertical offset of $G_{\mu\mu}(\tau)$ is arbitrary, and thus, we treat it as zero, making $\tilde{G}_{\mu\mu}(\tau)<0$ for all $\tau$. Again, the static Matsubara frequency $\omega_{n}=0$ must necessarily come out negative:
$G_{\mu\mu}(i\omega_{n}=0)=\int_{0}^{\beta}d\tau\tilde{G}_{\mu\mu}(\tau)<0,$ (36)
which stands out from the the rest, $G_{\mu\mu}(i\omega_{n}>0)>0$. Therefore, for convenience we will treat it with the opposite sign in figures and analyses, namely $\tilde{G}_{\mu\mu}(i\omega_{n}=0)\equiv-G_{\mu\mu}(i\omega_{n}=0)$. According to Eq. (11), $G_{\mu\mu}$ is directly connected to the physical polarizability for finite frequencies, i.e., $\alpha(\omega\neq 0)$. However, we will not seek here the ill-posed analytic continuation from imaginary to real domain.
Rather, for validation of the results for $\tilde{G}_{\mu\mu}$, we compare the PIMC data to the Drude model, given in Eq. (22), in the imaginary domain. Thus, we report the relative error $\Delta\tilde{G}_{\mu\mu}=(G_{\mu\mu}-G^{D}_{\mu\mu})/G^{D}_{\mu\mu}$ for both the corresponding autocorrelation and Matsubara functions. Unless stated otherwise, we use $\Gamma=0.0001$ in the Drude model to approximate low-$\Gamma$ limit while maintaining numerical stability. As will be seen, the correspondence is close and $\Delta\tilde{G}_{\mu\mu}\neq 0$ is rarely statistically significant. Therefore, for the first-order proxy of accuracy we will assume
$\eta\equiv\Delta\tilde{G}_{\mu\mu}(i\omega_{n}=0),$ (37)
which captures principal scaling differences between the PIMC and the Drude model, if any should be found.
Figure 2: Scaling of per-particle kinetic energy $K/N$ (left) and potential energy $V/N$ (bottom) with finite NN and variable dimensional Ewald cutoff parameter $r_{c}k_{c}$. The solid lines correspond to raw PIMC observables and the dotted lines are finite-size corrected values based on Eqs. (38) and (39). The case $r_{c}k_{c}=0$ means no analytical Ewald summation of any kind, including neutralizing background, and so the $E_{p}$ is off above the plot.
Let us considering systematic merits of accuracy for the PIMC simulation. Finite-size effects of the energies are studied by varying the number of particles $N$ in the simulation, while keeping the density fixed. Results for kinetic and potential energies are presented in Fig. 2 for up to $N=16$ and $\Theta=0.2$ at $r_{s}=4$. Shown are also finite-size corrections [40, 34] ($\xi=0$), which rectify the energies toward the thermodynamic limit (TDL), according to:
$K_{p}$ $=\frac{K_{N}}{N}+\frac{1}{N}\left(\frac{\omega_{p}}{4}+\frac{5.264}{4\pi r_{s}^{2}(2N)^{1/3}}\right)$ (38)
$V_{p}$ $=\frac{V_{N}}{N}+\frac{\omega_{p}}{4N},$ (39)
where the corrected energies $K_{p}$ and $V_{p}$ are further multiplied with $\mathrm{tanh(\omega_{p}\beta)}$. The correlation energy is getting saturated with $N$, which also influences the kinetic energy. Simulation with $N>16$, or a reliable extrapolation to the TDL, would be needed to resolve the statistically significant finite-size bias of the energies. Also, clearly the Ewald summation with $k_{c}r_{c}>10$ is required to eliminate statistically significant bias from treating periodic images.
Figure 3: Scaling of $\tilde{G}_{\mu\mu}(\tau)$ (left) and $\tilde{G}_{\mu\mu}(i\omega_{n})$ (right) with finite NN. The top panels show absolute values of the PIMC data (solid) and the corresponding Drude value (dotted). The bottom panels shows the relative differences $\Delta\tilde{G}_{\mu\mu}(\tau)$ (solid lines) and $\Delta\tilde{G}_{\mu\mu}(i\omega_{n})$, including statistical uncertainties (dotted lines).
To make physical benchmarks, there is obvious systematic bias related to the quantum statistics, i.e., the lack of Fermion exchange in the simulation. Otherwise, the Coulomb pair-density matrix is numerically exact and there are no nodal constraints. Remaining bias ensues from using a finite time-step $\Delta\tau$ to treat higher-body correlations (beyond pair approximation) and periodic images. As the higher-body correlations intensify with density, we assume for this work a choice of the time-step according to $\Delta\tau=ars^{3}$, where $a=0.001$ Ha-1.
The finite-size effects of the dipole polarizability are more topical: The scaling of $\tilde{G}_{\mu\mu}(\tau)$ and $\tilde{G}_{\mu\mu}(i\omega_{n})$ versus $N$ is plotted in Fig. 3 for $k_{c}r_{c}=15.0$. Overall, the effects of $\tilde{G}_{\mu\mu}$ from PIMC versus Drude due to finite size are negligible regardless of $N$. Nevertheless, we will use $N=16$ for the remainder of this work. The effects due to finite time-step and the Ewald cutoff parameter $k_{c}r_{c}$ are also statistically insignificant.
The lack of finite-size effects in $\tilde{G}_{\mu\mu}$ is in stark contrast with the energies, which are affected by the correlation and, although absent here, exchange. The indifference to the number of particles is therefore a manifestation of a greater principle: the trajectories of an ensemble of free particles in a periodic setting sample the phase space in collective harmony. The dipole autocorrelation estimator only measures the collective state and distributes it evenly for each particle. Another way of seeing this is that the low-momentum states of the plasma are occupied and cannot be used to excite the dielectric response of individual particles. Curiously, the cases of interacting (PIMC) and non-interacting (Drude) come out effectively the same.
Figure 4: On the left, $\tilde{G}_{\mu\mu}(\tau)$ and $\Delta\tilde{G}_{\mu\mu}(\tau)$ with variable $\Theta$ at $r_{s}=4$. On the right, $\tilde{G}_{\mu\mu}(i\omega_{n})$ and $\Delta\tilde{G}_{\mu\mu}(i\omega_{n})$ with variable $\Theta$ at $r_{s}=4$. Solid lines represent the absolute data and dotted lines the statistical uncertainties.
Finally, we study how $\tilde{G}_{\mu\mu}(\tau)$ depends on physical parameters, the finite temperature and density. In Fig. 4 we plot $\tilde{G}_{\mu\mu}(\tau)$ and $\tilde{G}_{\mu\mu}(i\omega_{n})$ versus $\Theta$ with $r_{s}=4$ and $N=16$. The plotting axes of $\tilde{G}_{\mu\mu}(\tau)$ are rescaled to $\tau/\beta$, emphasizing that the absolute values of $\tilde{G}_{\mu\mu}(\tau=\beta/2)$ decrease with the inverse temperature. Consequently, the Matsubara integral $\tilde{G}(i\omega_{n}=0)$ goes down with $\beta$, but otherwise the Matsubara series $\tilde{G}(i\omega_{n}>0)$ follow the same curve, only with different spacings. Clearly, the PIMC results match the Drude model within statistical uncertainty of $2\%$ or less.
In Figure 5 we discuss how the match depends on the Drude relaxation parameter $\Gamma$. When $\Gamma$ is increased, the Drude correlation function $\tilde{G}_{\mu\mu}(\tau)$ loses magnitude. The phenomenological explanation is that the scattering rate of the charge carriers from ions increases with $\Gamma$, and thus, the spatial diffusion slows down. The PIMC simulation of free particles has no scattering, and thus, the correspondence is the best with low $\Gamma$. This is presented in the lower panel of Fig. 5, where linear fits approach $\eta\rightarrow 0$ when $\Gamma\rightarrow 0$. The relative error $\eta$ scales linearly with $\Gamma$ and $\beta$, and so, curves representing $\eta/\beta$ are overlaid within statistical uncertainties. With this consideration we want to suggest a future possibility of estimating $\Gamma$ from PIMC simulations, provided a reasonable but yet unknown sampling method of the scattering process. As seen from all the presented results, such a simulation could be easily validated, while providing first-principles insights to the Drude scattering. Furthermore, using a finite-$\Gamma$ simulation such as that could also lend itself for studying nonlinear optical phenomena of Drude-like materials on a quantum mechanical level.
Figure 5: On the left, a comparison of $\tilde{G}_{\mu\mu}(\tau)$ versus $\tau$ from PIMC (solid black line) and the Drude model with variable values for $\Gamma$. On the right, the rescaled goodness measure $\eta/\beta$ is plotted against $\Gamma$ based on variables temperatures and $r_{s}=4$. Dotted lines are linear fits to the data.
V Summary and outlook
We have investigated a PIMC estimator for the dipole polarizability of periodic quantum gases and plasmas in the long-wavelength limit, i.e., the optical regime. We have bench marked the estimator with quantum Coulomb plasma in a range of finite densities ($r_{s}=2\ldots 8$ a0) and finite temperatures ($\Theta=0.1\ldots 0.8$ and up to $T=10^{6}$ Kelvin). The PIMC results are compared to the Drude model in the imaginary domain, where they match within statistical limits and $\Delta\tilde{G}_{\mu\mu}<2\%$, when the Drude damping coefficient $\Gamma$ is small. That is, the PIMC results for $\tilde{G}_{\mu\mu}$ show no statistically significant effects due (reasonable) simulation parameters, such as finite time-step, periodic images or finite number of particles, nor physical parameters such as finite density and temperature. This corroborates that the Drude spectrum is suitable for describing the linear optical response (i.e. the long-wavelength limit) regardless of the coupling strength. Perhaps some minor methodological, systematic biases would have surfaced with generous statistical sampling, but this does not appear as a worthy goal.
In summary: due to the collective motion of many interacting particles in a flat potential, all results come out consistent, effectively, with a free particle. This physical conclusion may not be overly interesting, but it firmly establishes that the numerical method, including the complete post-processing and analysis framework, is valid and robust. To this end, we have demonstrated special sampling approaches (see Sec. II.4), to allow accurate simulation even with a low particle count. These practices form a baseline for numerous further developments:
i) The Fermion spectrum. The treatment of signed permutations should be considered in the future, to unfold the true electron gas behaviour, including complexities in the statistical performance, sampling acceleration and path winding.
ii) Confined spectrum. Treating interacting particles in environments less regular than a flat potential (e.g. weak confinements, optical cavities or ionic plasma) should introduce non-trivial thermal and quantum effects for example with respect to the Drude–Lorentz model.
iii) Inverse Laplace transform. Special methods to invert Eqs. (16) or (17) could be studied to migrate possible quantum corrections into effective models. The MaxEnt method is not the best suited for the broad, Drude-like spectra [23, 41] of this work, but in the presence of resonance peaks, it could be.
iv) Higher-order response properties. The response functions beyond the dipole order could be more sensitive to thermal quantum many-body effects. Their static and dynamic variants, such as $\chi^{(3)}$ of the Drude-like materials, could inform effective models [5] and applications, such as ENZ materials [42].
Acknowledgements.
The authors wish to acknowledge CSC – IT Center for Science, Finland, and the Tampere Center for Scientific Computing; for computational resources. The authors acknowledge the financial support from the Photonics Research and Innovation Flagship (PREIN - decision 320165) and the Research Council of Finland project AQUA-PHOT (decision Grant No. 349350). JT acknowledges Ilkka Kylänpää for his support in the implementation and assessment of periodic potentials.

### v2

@@
 We present a path-integral Monte Carlo estimator for calculating the dipole polarizability of interacting Coulomb plasma in the long-wavelength limit, i.e., the optical region.
-Unlike the conventional dynamic structure factor in reciprocal space, our approach is based on the real-space dipole autocorrelation function and is suited for long wavelengths and small cell sizes, including finite clusters.
+We present comprehensive details and method validation studies for our approach based on dipole imaginary-time autocorrelation functions.
 The simulation of thermal equilibrium in imaginary time has exact Coulomb interactions and Boltzmann quantum statistics.
-For reference, we demonstrate analytic continuation of the Drude model into the imaginary time and Matsubara series, showing perfect agreement with our data within ranges of finite temperatures and densities.
-Method parameters, such as the finite time-step and finite-size effects prove only modestly significant.
-Our method, here carefully validated against an exactly solvable reference, remains amenable to more interesting domains in higher-order optical response, quantum confinements and quantum statistical effects, and applications in plasmonics, heterogeneous plasmas and nonlinear optics, such as epsilon-near-zero materials.
+For reference, we use the Drude model as the long-wavelength limit of the Lindhard response, presenting its analytical continuation into the imaginary time and Matsubara series.
+We demonstrate great agreement within $1\%$ between PIMC and the reference model, indicating negligible numerical biases and physical many-body effects in metallic densities.
+The approach is amenable to nonlinear optical response, quantum confinements, dispersion forces, and to inform applications such as plasmonics and epsilon-near-zero materials.
 preprint: APS/123-QED
@@
 This allows to characterize the behaviour of different photonic materials, such as metals [6, 7], plasmonic circuits [8], ENZ materials [9], as well as more complicated photonics platforms, such as metasurfaces [10], with an effective Drude model.
-Thus, there is great intrigue towards the recovery of the effective response properties from first principles.
+Yet, the Drude model remains phenomenological and unable to predict all subtleties of quantum mechanics from first principles.
 In spirit, the Drude model describes the dynamics of non-interacting free electrons, which occasionally scatter from ions.
-When the scattering is rare, this model becomes the homogeneous electron gas (HEG), or jellium, whose quantum mechanical linear response function is the Lindhard function $\chi_{0}(\mathbf{q},\omega))$ [11], where $\mathbf{q}$ and $\omega$ treat, respectively, the momentum and the frequency of the incoming field.
-Not surprisingly, the Lindhard function, despite begin famously hard to evaluate generally, recovers the Drude model upon taking approximations of long wavelength, high temperature and low scattering [7, 6].
-The Lindhard response can be expressed more generally as [12, 13]
-$\chi(\mathbf{q},\omega)=\frac{\chi_{0}(\mathbf{q},\omega)}{1+v_{q}(1-G(\mathbf{q},\omega))\chi_{0}(\mathbf{q},\omega)},$ (1)
-where the fraction treats dielectric screening, $v_{q}=e^{2}/q^{2}\varepsilon_{0}$ is the reciprocal Coulomb potential (in atomic units) and $G(\mathbf{q},w)$ is the so-called local field correction, basically treating effects due to Fermion exchange and dynamic correlations.
+When the scattering is modest, this model becomes the homogeneous electron gas (HEG), or jellium.
+Perhaps the most established framework for treating the dielectric response of HEG is the Lindhard theory, which builds upon the Lindhard function $\chi_{0}(\mathbf{q},\omega)$ [11], where $\mathbf{q}$ and $\omega$ are the respective wave number and frequency of an incoming perturbation.
+While somewhat tedious to evaluate numerically, $\chi_{0}$ recovers many physical insights in the right limits.
+In particular, the Drude model is recovered by taking $\chi_{0}$ to the limits of long wavelength and low scattering [7, 6].
+The Lindhard function cannot also describe the full many-body interactions of HEG, unless the Lindhard response is expressed more generally, like [12, 11]
+$\chi(\mathbf{q},\omega)$ $=-v_{q}\chi_{nn}(\mathbf{q},\omega)$ (1)
+$\chi_{nn}(\mathbf{q},\omega)$ $=\frac{\chi_{0}(\mathbf{q},\omega)}{1+v_{q}G(\mathbf{q},\omega)\chi_{0}(\mathbf{q},\omega)},$ (2)
+where $\chi$ denotes the complex dielectric susceptibility, $\chi_{nn}$ is the screened density–density susceptibility, $v_{q}=e^{2}/|\mathbf{q}|^{2}\varepsilon_{0}$ is the reciprocal Coulomb potential (in SI units) and $G(\mathbf{q},\omega)$ is the so-called local field correction treating many-body effects due to exchange and correlation.
+When $G=0$, Eq.
+(2) reduces to $\chi_{0}(\mathbf{q},\omega)$ and amounts to the random phase approximation.
 For decades, numerous works (e.g.
 Refs.
-[12, 14, 13]) have discussed and featured either static $G(\mathbf{q},\omega=0)$ or dynamic $G(\mathbf{q},\omega)$ in pursuit of the non-trivial response within the so-called regime of warm, dense matter (WDM).
-Recently, significant progress has been made to calculate the dielectric response of the WDM with the path-integral Monte Carlo method (PIMC) [15, 16, 17, 18, 19, 13, 20, 21].
-The PIMC approach can treat accurate quantum many-body interactions at non-trivial temperatures, and it is in principle only subject to numerical challenges due to the Fermion sign problem (FSP) [22, 23].
+[12, 13, 14, 15]) have discussed and featured either static $G(\mathbf{q},\omega=0)$ or dynamic $G(\mathbf{q},\omega)$ in pursuit of the non-trivial response within the so-called regime of warm, dense matter (WDM).
+Recently, significant progress has been made to calculate the density response of the WDM with the path-integral Monte Carlo method (PIMC) [16, 17, 18, 19, 20, 15, 21, 22].
+The PIMC approach can treat accurate quantum many-body interactions at finite temperatures, and it is in principle only subject to numerical challenges due to the Fermion sign problem (FSP) [23, 24].
 Within this framework, moreover, the dielectric response is directly connected to the dynamic structure factor through the intermediate scattering function [11].
@@
 have comprehensively studied the first [15] and higher order [18, 19] response properties of the HEG.
-An apparent limitation of their approach is that the allowed values for the wavenumber are limited by $q=2\pi n/L$, where $n\geq 1$ and $L$ is the (cubic) simulation cell size.
+An apparent limitation of their approach is that the allowed values for the wavenumber are limited by $|\mathbf{q}|=2\pi n/L$, where $n\geq 1$ and $L$ is the (cubic) simulation cell size.
 Therefore, the typical optical wavelengths with $q\ll L^{-1}$, thus requiring large $L$, are hard to accomplish because of the algorithmic scaling and performance loss attributed to the high particle count $N$.
@@
 Here, we explore a complementary PIMC approach to regain focus on the long, optical wavelengths.
-Namely, we demonstrate computation of the conventional optical response based on the complex dipole polarizability, or the dipole-dipole correlation function.
-This approach, which evades the difficulty of low $q$, has been studied before in finite systems [26, 27, 28, 29] but not in periodic ones.
-Granted, formal problems arise that are more naturally addressed in the Fourier space: The long-wavelength limit is only quasi-accurate, since we treat optical fields that are locally constant inside a finite box, i.e.
-$\lambda\gg L$, and formally expand that by summing over infinite periodic images.
-In practice, there is no field in the simulation, and thus, the approximation remains justified in all parameter ranges in the intended scope: for example, the box size $L$ required for $N=16$ particles to simulate the metallic density of gold is over $300$ smaller than an ultraviolet wavelength $\lambda=200$ nm.
-However, it is not obvious, a priori, how a PIMC estimator for the dipole polarizability must be laid out for periodic systems or how it responds to finite-size effects or converge parameters, such as the finite time-step.
-Therefore, it is the aim of this work to establish these matters in a controllable setting: We first discuss analytical correspondence between the Drude model and the PIMC polarizability, evaluated in the imaginary time.
-For convenience and consistency with the Drude model, we use distinguishable particles, i.e.
-the Boltzmann statistics, to evade the Fermion sign problem—for now.
+We, in fact, demonstrate computation of the conventional optical response based on the complex dipole polarizability, or the dipole-dipole correlation function.
+This approach, which evades the difficulty of low $\mathbf{q}$, has been studied before in finite systems [27, 28, 29, 30] but not in periodic ones.
+The long-wavelength approximation, where we treat a constant field in space, is well justified for optical region: For example, the box size $L$ required for $N=16$ particles to simulate the metallic density of gold is over $300$ times smaller than an ultraviolet wavelength $\lambda=200$ nm.
+Treating this naïvely in the Fourier space would require $N\sim 10^{8}$ particles in a cubic box.
+However, it is not obvious a priori how a PIMC estimator for the dipole polarizability must be laid out for periodic systems or how it responds to finite-size effects or converge parameters, such as the finite time-step.
+For these reasons, this work aims at establishing these matters in a controllable setting.
+To do so, we first discuss analytical correspondence between the Drude model and the PIMC polarizability, evaluated in the imaginary time.
+For convenience and consistency with the Drude model, we use distinguishable particles, i.e., Boltzmann statistics, to evade the Fermion sign problem, leaving a comprehensive investigation of its effects on our model to future works.
 We then survey various methodological and physical parameters to assert that the correspondence holds, when the Drude scattering rate is low.
-Indeed, we show that biases are statistically non-distinguishable within $2\%$ accuracy; the PIMC polarizability matches the Drude model.
-Formally, this is well anticipated, but practically the lack of biases is a signal of a robust simulation method.
-The assessments of performance help us avoid pitfalls, establish good practices and assert validity for continuation studies.
+Indeed, we show that biases are statistically non-distinguishable within $1\%$ relative accuracy; the PIMC polarizability matches the Drude model at metallic densities.
+While the result is not overly interesting in this regime, it asserts that the numerical method is robust and accurate for numerous continuation prospects.
 This work is organised as follows: In Sec.
@@
 In this section, we first overview the linear optical response of quantum particles in the long-wavelength limit.
-We then consider analytic continuation of the Drude model into imaginary domain, where it meets the dipole moment autocorrelation function accessible to the PIMC method.
-Throughout this work, we use the Hartree atomic units: $\hbar=1/(4\pi\epsilon_{0})=m_{e}=1$.
+We then discuss estimation of the dipole–dipole correlation function from imaginary-time PIMC simulation.
+Finally, we consider the Drude model as a non-interacting reference theory and its analytic continuation into the imaginary domain.
+The results of this section are presented in SI units, but the reader should keep in mind that later results will be given in in Hartree atomic units, i.e., $\hbar=4\pi\varepsilon_{0}=m_{e}=1$, for convenience.
 II.1 Electric field response
-The first-order local polarization $\mathbf{P}(\mathbf{r},\omega)$ of charged plasma responding to an incident electric field $\mathbf{F}(\mathbf{r^{\prime}},\omega)$ can be written as
-$\mathbf{P}(\mathbf{r},\omega)=\varepsilon_{0}\chi(\mathbf{r},\mathbf{r^{\prime}},\omega)\mathbf{F}(\mathbf{r^{\prime}},\omega),$ (2)
-where $\chi(\mathbf{r},\mathbf{r^{\prime}},\omega)$ is a complex linear response function, the electric susceptibility, in the local response approximation.
-Let us consider the long-wavelength limit, in which the field is oscillating with a wavelength $\lambda=2\pi c/\omega\gg L$, where $c$ is the speed of light and $L$ is a measure of longitudinal distance containing the interaction.
-Then, within $L$ we may consider a locally constant field
-$\mathbf{F}(\mathbf{r^{\prime}},\omega)\rightarrow\delta(\mathbf{r^{\prime}}-\mathbf{r})\mathbf{F}(\omega),$ (3)
-where $\delta$ is the Dirac delta function, inducing total mean polarization $P(\omega)\equiv\varepsilon_{0}\chi(\omega)\mathbf{F}(\omega)$ according to total mean susceptibility
-$\chi(\omega)=\frac{1}{L^{d}}\int_{0}^{L}\chi(\mathbf{r},\mathbf{r^{\prime}},\omega)\delta(\mathbf{r}-\mathbf{r^{\prime}})d\mathbf{r^{\prime}},$ (4)
-where $d$ is the dimensionality of $\mathbf{r}$. We may further consider the per-particle dipole polarizability $\alpha(\omega)$ through
-$\varepsilon_{0}\chi(\omega)=n\alpha(\omega),$ (5)
-where $n=N/L^{3}$ is the number density of the particles within $L$.
-The tensorial characters of $\chi$ and $\alpha$ are omitted here for simplicity of notation; because of the radial symmetry of the plasma, we only consider the diagonal, where $\mathbf{P}$ and $\mathbf{F}$ are parallel.
-For quantized plasma represented by $N$ particles, the per-particle dipole polarizability $\alpha(\omega)$ can also be expressed using the Kubo formula [30]
-$\alpha(\omega)$ $=-\frac{i}{N\hbar}\int_{0}^{\infty}\mathrm{d}te^{i\omega t}\int d\mathbf{r}\langle[\mu(\mathbf{r},t),\mu(\mathbf{r},0)]\rangle$ (6)
-$\equiv-\int_{-\infty}^{\infty}\mathrm{d}te^{i\omega t}G^{R}_{\mu\mu}(t)$ (7)
-$\equiv-G^{R}_{\mu\mu}(\omega),$ (8)
-where $G^{R}$ denotes the retarded dipole-dipole correlation function and the dipole moment operator is given by
-$\mu(\mathbf{r},t)=\sum_{i=1}^{N}q_{i}[\delta(\mathbf{r}-\mathbf{r_{i}}(t))-\langle\mathbf{r_{i}}\rangle],$ (9)
-where $q_{i}$ is the charge and $\mathbf{r_{i}}(t)$ the time-dependent displacement of the $i$th particle, and $\langle\mathbf{r_{i}}\rangle$ is their mean displacement, a non-perturbed reference point.
-Clearly, when $\sum q_{i}\neq 0$ the value depends on the reference point, which is problematic in translationally invariant systems, such as plasmas.
-Therefore, it is more meaningful to consider the relative correlation function
-$\tilde{G}^{R}_{\mu\mu}(t)\equiv G^{R}_{\mu\mu}(t)-G^{R}_{\mu\mu}(0)$ (10)
-which is zero at $t=0$ but nevertheless yields
-$\tilde{\alpha}(\omega)$ $\equiv-\int_{-\infty}^{\infty}\mathrm{d}te^{i\omega t}\tilde{G}^{R}_{\mu\mu}(t)=-G^{R}_{\mu\mu}(\omega)=\alpha(\omega)$ (11)
-for all $\omega\neq 0$, while $\alpha(\omega=0)$ remains elusive.
-On the other hand, it is well known that the total response of charged gas in the limits of $\omega=0$ and $q=0$ (long wavelength) is zero due to perfect screening [31].
-The linear response framework considered here does not treat screening, but this should justify why $\alpha(\omega=0)$ is unimportant for the free gas.
-The relative polarizability $\tilde{\alpha}(\omega)$ is analytic in the upper complex plane and can be analytically continued to imaginary time $it\rightarrow\tau$.
-The associated, time-ordered imaginary-time correlation function is given by
-$G_{\mu\mu}(\tau)=\langle\mathcal{T}\mu(0)\mu(\tau)\rangle,$ (12)
-where $\mathcal{T}$ is the time-ordering operator. The analytic continuation can be done via the spectral density function
-$A(\omega)\equiv-2\mathrm{Im}[G^{R}_{\mu\mu}(\omega)]$ (13)
-where the bosonic integral transformation yields
-$G_{\mu\mu}(\tau)=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\frac{\mathrm{e}^{-\tau\omega}}{1-\mathrm{e}^{-\beta\omega}}A(\omega),$ (14)
-where $G_{\mu\mu}(\tau)=G_{\mu\mu}(\tau+n\beta)$ for any integer $n$ and $\beta=1/k_{B}T$ is the inverse temperature. It is easy to see that
-$\tilde{G}_{\mu\mu}(\tau)$ $\equiv G_{\mu\mu}^{R}(\tau)-G_{\mu\mu}^{R}(0)$ (15)
-$=-\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\frac{1-\mathrm{e}^{-\tau\omega}}{1-\mathrm{e}^{-\beta\omega}}A(\omega).$ (16)
-Alternatively, we may write
-$G_{\mu\mu}(i\omega_{n})=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\frac{1}{i\omega_{n}-\omega}A(\omega),$ (17)
-where $\omega_{n}=2\pi n/\hbar\beta$ are the bosonic Matsubara frequencies.
-While the imaginary-domain functions $G_{\mu\mu}(\tau)$ and $G_{\mu\mu}(i\omega_{n})$ contain the same information as $G^{R}_{\mu\mu}(\omega)$, the conversion from imaginary to real domain by inverting Eqs.
+Let us consider a quantum mechanical plasma that is perturbed by a weak external electric field $\mathbf{E}(\mathbf{r^{\prime}},\omega)$ such that the Hamiltonian becomes
+$H\rightarrow H_{0}+H_{E}(\omega),$ (3)
+where $H_{0}$ is the non-perturbed Hamiltonian and $H_{E}(\omega)$ is a field-induced energy shift, integrated over volume $V$, is
+$H_{E}(\omega)$ $=\int_{V}d\mathbf{r}U_{E}(\mathbf{r})$ (4)
+$U_{E}(\mathbf{r})$ $=\mathbf{P}(\mathbf{r},\omega)\cdot\mathbf{E}(\mathbf{r},\omega),$ (5)
+where the induced electric polarization is prescribed to be linear with the (weak) field:
+$\mathbf{P}(\mathbf{r},\omega)\cong\varepsilon_{0}\int_{V}d\mathbf{r}^{\prime}\chi(\mathbf{r},\mathbf{r}^{\prime},\omega)\mathbf{E}(\mathbf{r^{\prime}},\omega),$ (6)
+where $\varepsilon_{0}$ the vacuum permittivity and $\chi(\mathbf{r},\mathbf{r}^{\prime},\omega)$ is the dielectric susceptibility tensor.
+By assuming translation invariance of plasma and regularity of $\mathbf{E}(\mathbf{r})$, we may write $\chi(\mathbf{r},\mathbf{r}^{\prime},\omega)=\chi(\mathbf{r}-\mathbf{r}^{\prime},\omega)$ and consider its Fourier transform $\chi(\mathbf{q},\omega)$ like in Eq.
+(1).
+On the other hand, we can make a long-wavelength approximation in the locality $L\sim V^{-3}$ such that $L\sim|\mathbf{r}-\mathbf{r^{\prime}}|\ll\lambda$, where $\lambda=2\pi/|\mathbf{q}|$ is the wavelength of the locally constant field $\mathbf{E}(\mathbf{r}-\mathbf{r}^{\prime},\omega)\cong\mathbf{E}(\omega)$.
+Thus, we can re-express Eq.
+(5) as a uniform potential energy density with
+$\delta U_{E}(\omega)$ $\cong\frac{1}{V}\int_{V}d(\mathbf{r}-\mathbf{r^{\prime}})\mathbf{P}(\mathbf{r}-\mathbf{r^{\prime}},\omega)\cdot\mathbf{E}(\omega)$ (7)
+$=\varepsilon_{0}|\mathbf{E}(\omega)|^{2}\chi(\omega)$ (8)
+where $\chi(\omega)\equiv V^{-1}\int_{V}d(\mathbf{r}-\mathbf{r^{\prime}})\chi(\mathbf{r}-\mathbf{r^{\prime}},\omega)$ is the bulk dielectric susceptibility, describing linear optical dispersion in the long-wavelength limit through the canonical relations [31]
+$\varepsilon(\omega)$ $=1+\chi(\omega)$ (9)
+$\mathbf{D}(\omega)$ $=\varepsilon_{0}\varepsilon(\omega)\mathbf{E}(\omega),$ (10)
+where $\varepsilon$ is the complex dielectric function and $\mathbf{D}(\omega)$ is the total electric displacement field.
+II.2 Quantum response functions
+Let us consider charged plasma comprising of $N$ identical quantum particles with charge e-e in a cubic box of volume $V=L^{3}$.
+For convenience, and without loss of generality, we can reduce the rank-2 susceptibility tensor $\chi(\omega)$ to its spherically symmetric diagonal, denoted $z$, and consider the response of this (homogeneous, isotropic) system to a long-wavelength electric field $E_{z}(\omega)$ applied along the direction $z$.
+From the zero-field limit of Eq.
+(8), we can derive
+$\varepsilon_{0}\chi(\omega)=\left.\frac{\partial^{2}}{\partial E_{z}^{2}}\frac{\delta U_{E}(\omega)}{V}\right|_{E_{z}\rightarrow 0},$ (11)
+which leads, using $U_{E}(\mathbf{r},\omega)=\mu(\mathbf{r})\cdot\mathbf{E}(\omega)$ and perturbation theory, to the Kubo formula [32]
+$\chi(\omega)$ $=\frac{n}{\varepsilon_{0}}\alpha(\omega)$ (12)
+$\alpha(\omega)$ $=-\frac{i}{\hbar}\int_{0}^{\infty}\mathrm{d}te^{i\omega t}\langle[\mu_{z}(t),\mu_{z}(0)]\rangle/N,$ (13)
+where $\alpha(\omega)$ is per-particle dipole polarizability, $n=N/V$ is the number density, the angle brackets evaluate a quantum statistical average over the volume $V$, and square brackets denote a commutator.
+The dipole moment operator in $z$ direction is generally
+$\mu_{z}(\mathbf{r},t)=\sum_{i=1}^{N}q_{i}z_{i}\delta(\mathbf{r}-\mathbf{r_{i}}(t)),$ (14)
+where $q_{i}$ and $\mathbf{r_{i}}(t)$ are the respective charge and position of the $i$th particle.
+If $|q_{i}|=e$ for all $i$, Eqs.
+(12) and (13) can be expressed with the Fourier-space density–density susceptibility $\chi_{nn}(\omega)$ like in Eq.
+(1) or the dynamic structure factor [11], as is typically done to treat shorter wavelengths.
+Let us denote the dipole–dipole correlation function with
+$\frac{i}{\hbar}\langle[\mu(t),\mu(0)]\rangle/N\equiv G(t)$ (15)
+such that the retarded (causal) correlation function is
+$G^{R}(\omega)=\int_{-\infty}^{\infty}\mathrm{d}te^{i\omega t}\Theta(t)G(t)=-\alpha(\omega),$ (16)
+whose spectral density function is [25]
+$A(\omega)\equiv-2\mathrm{Im}[G^{R}(\omega)].$ (17)
+The spectral density allows analytic continuation of $G^{R}$ to imaginary time $it\rightarrow\tau$ through [25]
+$\mathcal{G}(\tau)=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\frac{\mathrm{e}^{-\tau\omega}}{1-\mathrm{e}^{-\beta\omega}}A(\omega),$ (18)
+where $\beta=1/k_{B}T$ is the inverse temperature.
+The imaginary-time correlation function $\mathcal{G}(\tau)$ has bosonic symmetry with $\mathcal{G}(\tau)=\mathcal{G}(\tau+n\beta)$ for any integer $n$, and similar to its counterpart $G^{R}(t)$, is evaluated with
+$\mathcal{G}(\tau)=\frac{1}{\hbar}\langle\mathcal{T}\mu_{z}(0)\mu_{z}(\tau)\rangle/N,$ (19)
+where $\mathcal{T}$ is the time-ordering operator.
+The dipole moment operators $\mu_{z}(\tau)$ are also assumed normal ordered, meaning $\langle\mu_{Z}(\tau)\rangle=0$.
+The Fourier transform of $\mathcal{G}(\tau)$ is the Matsubara series $\mathcal{G}(i\omega_{n})$, where $\omega_{n}=2\pi n/\hbar\beta$ for all integers, which can also be obtained with [25]
+$\mathcal{G}(i\omega_{n})=-\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\frac{1}{i\omega_{n}-\omega}A(\omega).$ (20)
+While the imaginary-domain functions $\mathcal{G}(\tau)$ and $\mathcal{G}(i\omega_{n})$ contain the same information as $G^{R}(\omega)$, the conversion from imaginary to real domain by inverting Eqs.
 (18) or (20) is a known ill-posed problem, which is accentuated by the unavoidable error bars present in quantum Monte Carlo methods [25].
-The use of existing numerical methods of analytic continuation, such as Maximum Entropy (MaxEnt) [25, 26], will be left for future consideration.
-II.2 Analytic continuation of the Drude model
-The ubiquitous model for considering the optical response of microscopic particles is the Drude model. The Drude susceptibility is given by
-$\chi^{D}(\omega)=-\frac{\omega_{p}^{2}}{\omega^{2}+i\omega\Gamma},$ (18)
-where $\omega_{p}^{2}=\tfrac{ne^{2}}{m\varepsilon_{0}}=\tfrac{e^{2}}{4\pi\varepsilon_{0}m}\tfrac{3}{r_{s}^{3}}$ is the plasma frequency and $\Gamma$ is a positive damping coefficient.
-The Drude per-particle polarizability is
-$\alpha^{D}(\omega)=\frac{\chi_{e}^{D}(\omega)\varepsilon_{0}}{n},$ (19)
-whose spectral density, according to Eq (13), is given by
-$A^{D}(\omega)$ $=\frac{2\omega_{p}^{2}\varepsilon_{0}}{n}\frac{1}{\Gamma\omega(1+\omega^{2}/\Gamma^{2})}$ (20)
-$=\frac{2e^{2}}{m}\frac{1}{\Gamma\omega(1+\omega^{2}/\Gamma^{2})}.$ (21)
-The spectrum diverges at $\omega=0$ and its analytic continuation with Eq.
-(14) becomes problematic.
-However, the relative polarizability $\tilde{\alpha}(\omega)$ can be obtained from Eq.
-(16) because the divergence is controlled at $\omega=0$ by the transformation kernel.
-Furthermore, an analytical expression for $G^{D}_{\mu\mu}(i\omega_{n})$, which is the correlation function in Matsubara frequencies associated to the Drude model, can be obtained by using Eq.
-(20) and substituting into Eq.
-17.
-This yields
-$G^{D}_{\mu\mu}(i\omega_{n})=\frac{2\pi e^{2}}{m}\left[\frac{1}{\omega_{n}^{2}/\Gamma+\omega_{n}}\right].$ (22)
-The analytical expression given by Eq.  (22) is valid for $\omega_{n}>0$. The static limit is defined here based on the zero-frequency Fourier transform of relative correlation function
-$\tilde{G}^{D}_{\mu\mu}(i\omega_{n})=\int_{0}^{\beta}d\tau\tilde{G}^{D}_{\mu\mu}(\tau),$ (23)
-but as discussed before, it is not physical and only used for numerical assessment of performance.
-From Eq.
-(20) it is clear that the Drude spectrum does not depend on the physical density, because the oscillators are not interacting; rather, the density only shows in the transformation to the macroscopic susceptibility with Eq.(5).
-To consider the interacting problem, let us therefore turn to numerical path-integral simulation.
+The use of existing numerical methods of analytic continuation, such as Maximum Entropy [26, 27], will not be pursued here.
+Figure 1: Schematic illustration of the periodic dipole moment estimator and sampling constraints.
+Blue circles represent Monte Carlo walkers of quantum particles, i.e., positions advancing in discrete steps through imaginary time and forming closed loops $z(0)=z(\beta)$ modulo $L$, where $z$ denotes the real coordinate in one dimension.
+(left) The mean dipole autocorrelation function is easily measured by summing up individual particles, here represented by $\mu_{z}(\tau)/qN$ for $N=2$ by the smaller green dots.
+Their mean squared fluctuation w.r.t.
+to the mean (green vertical line) induces the relative autocorrelation function $\mathcal{\tilde{G}}(\tau)$ (visualized by the width of a green distribution).
+(middle) Unfolding of the path.
+Since the estimator does not treat periodicity with a Fourier phase, paths crossing the simulation box boundary to the nearest image must be unfolded, i.e., moved within $L/2$ of the earlier bead, before measurement.
+(right) Winding restriction.
+MC sampling may cause the path to wind across box boundary (branch of red dots), when the thermal wavelength of a multilevel bisection move is a significant fraction of $L$.
+For distinguishable particles, this is clearly an artifact of the finite cell and also detrimental to the estimation of $\mathcal{\tilde{G}}(\tau)$, and thus, such moves are rejected.
 II.3 PIMC estimator
-The PIMC method can be used to simulate interacting quantum particles at finite temperatures. The method draws samples from the distribution of the thermal density matrix:
+The quantum statistical averages of the dipole–dipole correlation function at finite temperatures can be obtained with the canonical PIMC method [24].
+The method draws samples from the distribution of the thermal density matrix:
 $\rho(R,R;\beta)=Z^{-1}\int dR\langle R|\hat{\rho}(\beta)|R\rangle,$ (21)
-where $R$ denotes a set of variable particle coordinates, $\hat{\rho}(\beta)=e^{-\beta H}$ is the density operator at inverse temperature $\beta$, $H$ is the Hamiltonian and $Z=\langle R|\hat{\rho}|R\rangle$ is the thermal partition function.
-Equation (24) overlooks summations arising from spin degrees of freedom and particle permutations required to treat proper quantum statistics of bosons or fermions; instead, the simulation treats so-called boltzmannons, which comprise a Coulomb interacting gas of distinguishable quantum particles.
+where $R$ denotes a set of variable particle coordinates, $\hat{\rho}(\beta)=e^{-\beta H_{0}}$ is the density operator at inverse temperature $\beta$, $H_{0}$ is the Hamiltonian and $Z=\langle R|\hat{\rho}|R\rangle$ is the thermal partition function.
+Equation (21) overlooks summations arising from spin degrees of freedom and particle permutations required to treat proper quantum statistics of bosons or fermions; instead, the simulation treats so-called Boltzmannons, which comprise a Coulomb interacting gas of distinguishable quantum particles [33].
 The Boltzmannons evade the FSP but do not correspond to physical electrons, alas, except at high temperatures and low densities [24] (or strong coupling [34]), where particle repulsion outweighs exchange and the quantum statistics merge.
-Our results are also affected by the (lack of proper) quantum statistics.
+As elaborated in the SI, some of our results indeed start to deviate from the Fermion values.
 The interacting Hamiltonian for $N$ distinguishable particles is given by
-$H(R)=T+V=\lambda_{m}\sum_{i}\nabla_{i}^{2}+\sum_{i<j}\frac{q_{i}q_{j}}{|\mathbf{r}_{i}-\mathbf{r}_{j}|}$ (25)
-where $\lambda_{m}=1/2m_{e}$ and $T$ and $V$ correspond, respectively, to the summation over kinetic and potential operators, the last ones comprising exact Coulomb interaction between pairs of particles.
-Since $T$ and $V$ do not commute, the density operator must be expanded with the semi-group property
-$e^{-\beta(T+V)}=\lim_{M\rightarrow\infty}\left[e^{-\Delta\tau T}e^{-\Delta\tau V}\right],$ (26)
-where $\Delta\tau=\beta/M$ is the time-step and $M$ is the Trotter number.
+$H_{0}=K+U=\hbar^{2}\lambda_{m}\sum_{i}\nabla_{i}^{2}+\sum_{i<j}\frac{q_{i}q_{j}}{4\pi\varepsilon_{0}|\mathbf{r}_{i}-\mathbf{r}_{j}|}$ (22)
+where $\lambda_{m}=1/2m_{e}$ and $K$ and $U$ correspond, respectively, to the summation over kinetic and potential operators, the last ones comprising exact Coulomb interaction between pairs of particles.
+Since $K$ and $U$ do not commute, the density operator must be expanded with the semi-group property
+$e^{-\beta(K+U)}=\lim_{M\rightarrow\infty}\left[e^{-\Delta\tau K}e^{-\Delta\tau U}\right],$ (23)
+where $\Delta\tau=\hbar\beta/M$ is a finite time-step and $M$ is the Trotter number.
 Using Eq.
@@
 The trajectory is efficiently sampled as a Markov chain using the Metropolis multilevel bisection algorithm, as described in [33, 24, 35].
-The calculation of the thermal expectation values of diagonal properties, such as potential energy $V$, is done from Eq. (27) as
-$\langle V\rangle$ $=Z^{-1}\int dR\langle R|\hat{V}\hat{\rho}(\beta)|R\rangle$ (28)
-$=Z^{-1}\prod_{j=1}^{M}dR_{j}\rho(R_{j-1},R_{j};\Delta\tau)V(R_{0})$ (29)
-$=\frac{1}{M}\sum_{j=1}^{M}\langle V(R_{j}),$ (30)
-where the last line employs symmetry properties of the trace, enabling to calculate mean over all time-slices of any trajectory.
+Thermal expectation values of diagonal properties, such as potential energy $V$, is can be calculated from
+$\langle U\rangle$ $=Z^{-1}\int dR\langle R|U\hat{\rho}(\beta)|R\rangle$ (25)
+$\approx Z^{-1}\int\left[\prod_{j=1}^{M}dR_{j}\rho\left(R_{j-1},R_{j};\Delta\tau\right)\right]U(R_{0})$ (26)
+$=\frac{1}{M}\sum_{j=1}^{M}\langle U(R_{j})\rangle,$ (27)
+where the last line employs symmetry properties of the trace, enabling to calculate the mean over all time-slices of any trajectory.
 Because of the kinetic operator, the total energy $E$ is more complicated; here we use the thermal estimator [24], whose numerically exact pair energy $\partial\rho/\partial\beta$ can be precomputed [35].
-The calculation of time-ordered observables, such as the imaginary-time correlation function from Eq. (12), is equally straightforward. The mean per-particle correlation function is evaluated as
-$\langle G_{\mu\mu}(\tau)\rangle=\frac{1}{ZN}\prod_{j=1}^{M}dR_{j}\rho(R_{j-1},R_{j};\Delta\tau)\mu(R_{0})\mu(R_{t}),$ (31)
-where the time-difference index is $t=j+\tau/\Delta\tau$, modulo $M$.
-The expression can be easily symmetrized, optimized for evaluation and generalized for different combinations of measurement operators, as detailed in [28, 29].
-In the path-integral context, the dipole moment operator is
-$\mu(R_{t})$ $=\sum_{i=1}^{N}q_{i}\left[\mu(\mathbf{r}_{i}(t\Delta\tau)-\langle\mu_{i}\rangle\right],$ (32)
-which again depends on the point of reference $\langle\mu_{i}\rangle$. However, as was done in Ref. [26], we identify
-$\langle\tilde{G}_{\mu\mu}(\tau)\rangle$ $=\frac{1}{2N}\langle(\mu(0)-\mu(\tau))^{2}\rangle,$ (33)
-where the relative correlation function arises from mean squared fluctuations of the instantaneous dipole moment along the imaginary-time path.
+The calculation of time-ordered observables such as Eq. (19) is equally straightforward. The mean per-particle correlation function can evaluated at discrete multiples of $\Delta\tau$ by [30]
+$\langle\mathcal{G}(\tau)\rangle=\frac{1}{\hbar ZN}\prod_{j=1}^{M}\int dR_{j}\rho(R_{j-1},R_{j};\Delta\tau)\mu_{z}(R_{0})\mu_{z}(R_{t}),$ (28)
+where the integer time-difference index is $t=j+\tau/\Delta\tau$, modulo $M$.
+A snapshot measurement of the correlation function from the trajectory is visualized in Fig.
+1 (left).
+The expression Eq.
+(28) can be easily symmetrized, optimized for evaluation and generalized for different combinations of measurement operators, as detailed in [29, 30].
+The imaginary-time dipole moment operating on a time-slice is
+$\mu_{z}(R_{t})$ $=\sum_{i=1}^{N}q_{i}\left(z_{t}-z_{0}\right),$ (29)
+where $z_{0}$ is a reference point for $z$.
+A difficulty arises for free plasmas, for which $\sum_{i}q_{i}\neq 0$ and there is no extrinsic reference point.
+One might be tempted to use an intrinsic reference point, such as the path-centroid $z_{0}\equiv M^{-1}\sum_{t}z_{t}$, but that leads to $\langle\mathcal{G}(i\omega_{n}=0)=0$.
+The lack of a reference point is a fundamental consequence of the translation invariance, gracefully treated in the reciprocal space.
+Therefore, we shall not directly evaluate $\langle\mathcal{G}(\tau)\rangle$ and instead (like in Ref. [27]) consider the relative correlation function
+$\mathcal{\tilde{G}}(\tau)$ $\equiv\mathcal{G}(\tau)-\mathcal{G}(0)$ (30)
+$=-\frac{1}{2\hbar}\langle(\mu(0)-\mu(\tau))^{2}\rangle/N,$ (31)
+which arises from mean squared fluctuations of the instantaneous dipole moment along the imaginary-time path (the green path in Fig.
+1 (left)).
+This is not such a huge compromise, as the relative correlation functions in real domain follow
+$\tilde{G}^{R}(t)$ $\equiv G^{R}(t)-G^{R}(0)$ (32)
+$\tilde{G}^{R}(\omega)$ $=\int dt\mathrm{e}^{i\omega t}\tilde{G}^{R}(t)=-\alpha(\omega\neq 0),$ (33)
+so we can, in principle, treat all values of $\alpha(\omega)$ except the static value.
 II.4 Periodic boundary conditions
-The simulation of charged plasma requires periodic boundary conditions (PBC). Here we use a cubic box of size $L$, whose size is determined by
-$L=r_{s}\left(\frac{4\pi N}{3}\right)^{(1/3)},$ (34)
-where $r_{s}=(3/4\pi n)^{1/3}$ is the Wigner-Seitz radius of the particle gas.
-The treatment of periodic Coulomb images is treated using the ordinary Ewald summation [36, 37].
-Note that the long-wavelength approximation discussed in Sec.
-II.1 also uses $L$, and the symbols are in practice interchangeable but semantically different.
-The finite density is enforced for the many-body interactions of the plasma, but otherwise the evaluation resembles an open-boundary problem.
-So, the evaluation of the dipole polarizability estimator, Eq.
-(33), is done on walkers whose all particle paths are unfolded into continuous trajectories near boundaries.
+The simulation of charged plasma requires periodic boundary conditions (PBC).
+Here we use a cubic box of size $L$ (in practice like in Sec.
+II.1 although semantically different), whose size is determined by the
+$L=r_{s}\left(\frac{4\pi N}{3}\right)^{1/3},$ (34)
+where $r_{s}$ is the Wigner-Seitz radius.
+The treatment of periodic Coulomb images is treated using the ordinary Ewald summation [37, 38] (see SI for details).
+Evaluation of the relative dipole autocorrelation function (Eqs.
+(28) and (30)), is done on walkers whose all particle paths are unfolded into continuous trajectories near boundaries, as illustrated in Fig.
+1 (middle).
 Since the simulated particles may not exchange paths (permute), non-zero winding of a path through the cell is an artifact of the finite size.
-Therefore, we reject all moves leading to non-zero winding.
+Therefore, we reject all moves leading to non-zero winding (see Fig.
+1 (right)).
 Beside this constraint, the sampling is done using the multilevel bisection method [24].
-The probability of winding increases when the thermal wavelength of the bisection move approaches the cell dimension, i.e.
+The probability of winding increases when the thermal de Broglie wavelength of the bisection move approaches the cell dimension, i.e.
 $\sqrt{2^{l}\lambda_{m}\Delta\tau}\sim L,$ (35)
@@
 This is a simplification to avoid using periodic free-particle density matrix [24, 35], which would not inhibit the winding, anyway.
-As a result, fewer than one winding move per $10^{3}$ trial moves is observed and has to be rejected in any density.
+As a result, fewer than one winding move per $10^{3}$ trial moves needs to be rejected in any density.
 A modest sampling bias is mitigated along other finite-size effects when $L$ naturally grows with $N^{1/3}$.
+II.5 Analytic continuation of the Drude model
+The ubiquitous model for considering the optical response free-carrier species like electrons in metals is the Drude model, whose dielectric susceptibility is
+$\varepsilon^{D}(\omega)=1-\frac{\omega_{p}^{2}}{\omega^{2}+i\omega\Gamma},$ (36)
+and thus, the susceptibility is
+$\chi^{D}(\omega)=-\frac{\omega_{p}^{2}}{\omega^{2}+i\omega\Gamma},$ (37)
+where $\omega_{p}^{2}=\tfrac{ne^{2}}{m\varepsilon_{0}}=\tfrac{e^{2}}{4\pi\varepsilon_{0}m}\tfrac{3}{r_{s}^{3}}$ is the plasma frequency and $\Gamma>0$ is a positive damping coefficient.
+As is well known [6] and iterated in Appendix A, the Drude model is recovered from the Lindhard function $\chi_{0}(\mathbf{q},\omega)$ in the limits of small $\Gamma$ and long wavelength as $q\rightarrow 0$.
+Thus, the Drude model is used in this work as a reference model that derives from quantum mechanics and contains screening on mean-field level [11] but neglects many-body effects from exchange and correlation.
+Differences between the PIMC data and Drude model are therefore manifestations of the many-body effects.
+Based on Eqs. (12) and (37), the Drude polarizability per particle is given by
+$\alpha^{D}(\omega)=\frac{\chi^{D}(\omega)\varepsilon_{0}}{n},$ (38)
+whose spectral density, according to Eq (17), is
+$A^{D}(\omega)$ $=\frac{2\omega_{p}^{2}\varepsilon_{0}}{n}\mathrm{Im}[\chi^{D}(\omega)]$ (39)
+$=\frac{2e^{2}}{m}\frac{1}{\omega(\Gamma+\omega^{2}/\Gamma)}.$ (40)
+From Eq.
+(39) it is clear that the Drude spectrum does not depend on the physical density, because the oscillators are not interacting; rather, the density only shows in the transformation to the macroscopic susceptibility with Eq.(12).
+The spectrum also diverges at $\omega=0$ complicating its analytic continuation into $\mathcal{G}$.
+However, the relative correlation function $\mathcal{\tilde{G}}$ can be obtained, because the subtraction of $G^{R}(\tau=0)$ controls the divergence at $\omega=0$.
+So, based on Eqs.(39), (18) and (20) we can perform numerical analytic continuation with
+$\mathcal{\tilde{G}}^{D}(\tau)$ $=-\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\frac{1-\mathrm{e}^{-\tau\omega}}{1-\mathrm{e}^{-\beta\omega}}A^{D}(\omega).$ (41)
+For the Matsubara data we get an analytic expression for the non-zero Matsubara frequencies:
+$\mathcal{\tilde{G}}^{D}(i\omega_{n}\neq 0)$ $=\mathcal{G}^{D}(i\omega_{n}\neq 0)$ (42)
+$=\frac{2\pi e^{2}}{m}\left[\frac{1}{\omega_{n}^{2}/\Gamma+\omega_{n}}\right],$ (43)
+However, we can formally define the static limit based on the zero-frequency Fourier transform
+$\mathcal{\tilde{G}}(i\omega_{n}=0)\equiv-\int_{0}^{\beta}d\tau\mathcal{\tilde{G}}(\tau),$ (44)
+where the sign has been changed for convenience, to keep all values of $\mathcal{\tilde{G}}(i\omega_{n})>0$ in figures.
+As discussed before, $\mathcal{\tilde{G}}(i\omega_{n}=0)$ is not physical but since it describes the first-order scaling of the correlation function, we will use it later as a measure of quantity.
 III Computational details
-The PIMC simulation software, including the exact Coulomb pair action generation, implements the canonical Metropolis Monte Carlo algorithm for a PIMC walker in Fortran90, featuring parallel sampling and data binning of a number of Markovian walkers.
-Here, 40 CPU cores are used for parallel sampling.
+The PIMC simulation software implements the canonical Metropolis Monte Carlo algorithm for a PIMC walker in Fortran90, featuring matrix squaring of the exact Coulomb pair action, parallel sampling and data binning of a number of Markovian walkers.
+Here, 40 CPU cores are used for the parallel sampling.
 Raw data after sample binning is stored in hierarchical data format (HDF5) and data processing chains operate on standard numerical Python libraries.
 The integral transformations, such as Eq.
+(18) and Eq.
 (20), are carried out using numerical quadrature.
-The simulation campaigns are operated with Nexus workflows [38] on various high-performance computing (HPC) facilities listed in Sec.
-V.
-All implementation details, including code snapshots and workflow scripts for step-by-step reproduction of the simulations, postprocessing and individual figures are supplied in a separate data repository [39].
+Fourier transforms from $\mathcal{\tilde{G}}(\tau)\rightarrow\mathcal{G}(i\omega_{n})$ are carried on regular grids of finite spacing using fast Fourier transform (FFT); however, FFT biases the result toward higher frequencies, and thus, the original data is interpolated 16-fold while truncating the size of the result array to mitigate the bias [26, 29].
+The simulation campaigns are operated with Nexus workflows [39] on various high-performance computing facilities listed in Sec.
+VI.
+Central implementation details, including code snapshots and workflow scripts for step-by-step reproduction of the simulations, postprocessing and individual figures are supplied in a separate data repository [40].
 […]
 IV Results
-Figure 1: Comparison overview of the relative dipole correlation function $\tilde{G}_{\mu\mu}$ from PIMC and the Drude model.
-In the top, the imaginary-time correlation function $\tilde{G}_{\mu\mu}(\tau)$ is plotted at two temperatures, $\Theta_{1}=0.1$ and $\Theta_{2}=0.2$, resulting in two curves ending at different values for $\beta_{1}=2\beta_{2}$.
-In the bottom, the first few corresponding Matsubara data $\tilde{G}_{\mu\mu}(i\omega_{n})$ are plotted, where the frequencies $\omega_{n}=2\pi n/\hbar\beta$ for $\Theta_{1}$ are twice those of $\Theta_{2}$, but the two functions lie on the same curve.
+The PIMC simulation, as laid out in earlier section, is applied in quantum plasmas at various finite temperatures and densities.
+With $m=m_{e}=1$ and $q^{2}=e^{2}=1$, the simulated particles resemble free electrons in the boltzmannon approximation.
+Thus, we use here conventional units of the HEG and express densities with $r_{s}$ and finite temperatures with $\Theta=T/T_{f}$, where $k_{B}T_{f}=\hbar^{2}k_{f}^{2}/2$ and $k_{f}=(3n\pi^{2})^{1/3}$ are, respectively, the Fermi temperature and the Fermi wavevector.
+The covered densities range between $r_{s}=2.0\ldots 12$ a0 and finite temperatures between $\Theta=T/T_{f}=0.1\ldots 0.8$.
+The main observables of interest are overviewed in Fig.
+2: the relative dipole autocorrelation function $\mathcal{\tilde{G}}(\tau)$ and its Matsubara series $\mathcal{\tilde{G}}(i\omega_{n})$.
+As argued in Sec.
+II, the vertical offset of $\mathcal{\tilde{G}}(\tau)$ is arbitrary, and thus, we treat it as zero, making $\mathcal{\tilde{G}}(\tau)\leq 0$ for all $\tau$.
+Thus, data for the static Matsubara frequency would be numerically negative unless changed defined with an opposite sign like in Eq.
+(44).
+In the absence of notable many-body effects, all values of $\mathcal{\tilde{G}}(i\omega_{n})$ lie on the same curve, except $\mathcal{\tilde{G}}(i\omega_{n}=0)$ which informs the first-order scale of the correlation function.
+Figure 2: Comparison overview of the relative dipole correlation function $\mathcal{\tilde{G}}$ from PIMC and the Drude model.
+In the top, the imaginary-time correlation function $\mathcal{\tilde{G}}(\tau)$ is plotted at two temperatures, $\Theta_{1}=0.1$ and $\Theta_{2}=0.2$, resulting in two curves ending at different values for $\beta_{1}=2\beta_{2}$.
+In the bottom, the first few corresponding Matsubara data $\mathcal{\tilde{G}}(i\omega_{n})$ are plotted, where the frequencies $\omega_{n}=2\pi n/\hbar\beta$ for $\Theta_{1}$ are twice those of $\Theta_{2}$, but the two functions lie on the same curve.
 In both cases, the data match the analytically continued Drude model within statistical uncertainties.
-Let us now discuss results of the PIMC simulation, laid out in the earlier sections, as applied to charged quantum plasma at various finite temperatures and densities.
-With $m=m_{e}$ and $q^{2}=e^{2}$, the simulated particles resemble free electrons, only without sign-flipping permutations.
-Thus, we use here conventional units of the HEG and express densities with $r_{s}$ and finite temperatures with $\Theta=T/T_{f}$, where $k_{B}T_{f}=\hbar^{2}k_{f}^{2}/2$ and $k_{f}=(3n\pi^{2})^{1/3}$ are, respectively, the Fermi temperature and the Fermi wavevector.
-The covered densities range between $r_{s}=2.0\ldots 8$ a0 and finite temperatures between $\Theta=T/T_{f}=0.1\ldots 0.8$.
-The main observables of interest are overviewed in Fig.
-1: the relative imaginary-time correlation function $\tilde{G}_{\mu\mu}(\tau)$ and its Matsubara series $G_{\mu\mu}(i\omega_{n})$.
-As stated in Sec.
-II, the vertical offset of $G_{\mu\mu}(\tau)$ is arbitrary, and thus, we treat it as zero, making $\tilde{G}_{\mu\mu}(\tau)<0$ for all $\tau$.
-Again, the static Matsubara frequency $\omega_{n}=0$ must necessarily come out negative:
-$G_{\mu\mu}(i\omega_{n}=0)=\int_{0}^{\beta}d\tau\tilde{G}_{\mu\mu}(\tau)<0,$ (36)
-which stands out from the the rest, $G_{\mu\mu}(i\omega_{n}>0)>0$.
-Therefore, for convenience we will treat it with the opposite sign in figures and analyses, namely $\tilde{G}_{\mu\mu}(i\omega_{n}=0)\equiv-G_{\mu\mu}(i\omega_{n}=0)$.
-According to Eq.
-(11), $G_{\mu\mu}$ is directly connected to the physical polarizability for finite frequencies, i.e., $\alpha(\omega\neq 0)$.
-However, we will not seek here the ill-posed analytic continuation from imaginary to real domain.
-Rather, for validation of the results for $\tilde{G}_{\mu\mu}$, we compare the PIMC data to the Drude model, given in Eq.
-(22), in the imaginary domain.
-Thus, we report the relative error $\Delta\tilde{G}_{\mu\mu}=(G_{\mu\mu}-G^{D}_{\mu\mu})/G^{D}_{\mu\mu}$ for both the corresponding autocorrelation and Matsubara functions.
-Unless stated otherwise, we use $\Gamma=0.0001$ in the Drude model to approximate low-$\Gamma$ limit while maintaining numerical stability.
-As will be seen, the correspondence is close and $\Delta\tilde{G}_{\mu\mu}\neq 0$ is rarely statistically significant.
-Therefore, for the first-order proxy of accuracy we will assume
-$\eta\equiv\Delta\tilde{G}_{\mu\mu}(i\omega_{n}=0),$ (37)
-which captures principal scaling differences between the PIMC and the Drude model, if any should be found.
-Figure 2: Scaling of per-particle kinetic energy $K/N$ (left) and potential energy $V/N$ (bottom) with finite NN and variable dimensional Ewald cutoff parameter $r_{c}k_{c}$.
+Figure 3: Scaling of per-particle kinetic energy $K$ (left) and potential energy $V$ (bottom) with finite $N$ and variable dimensional Ewald cutoff parameter $r_{c}k_{c}$.
 The solid lines correspond to raw PIMC observables and the dotted lines are finite-size corrected values based on Eqs.
 (47) and (48).
-The case $r_{c}k_{c}=0$ means no analytical Ewald summation of any kind, including neutralizing background, and so the $E_{p}$ is off above the plot.
-Let us considering systematic merits of accuracy for the PIMC simulation.
+The case $r_{c}k_{c}=0$ means no analytical Ewald summation of any kind, including neutralizing background, and so $V$ is off above the plot.
+We generally seek traces of methodological and physical discrepancies between the PIMC data and the Drude model. The differences are treated in terms of relative error
+$\Delta\mathcal{\tilde{G}}=(\mathcal{\tilde{G}}-\mathcal{\tilde{G}}^{D})/\mathcal{\tilde{G}}^{D}$ (45)
+for any imaginary arguments $\tau$ or $i\omega_{n}$.
+Unless stated otherwise, we use $\Gamma=0.0001$ in the Drude model to approximate low-$\Gamma$ limit while maintaining numerical stability.
+As will be seen, the correspondence is close and $\Delta\mathcal{\tilde{G}}\neq 0$ is rarely statistically significant.
+Therefore, for a first-order measure of accuracy we shall use
+$\eta\equiv\Delta\mathcal{\tilde{G}}(i\omega_{n}=0),$ (46)
+which captures principal scaling differences between the PIMC and the Drude model, if any should be found.
+Figure 4: Scaling of $\mathcal{\tilde{G}}(\tau)$ (left) and $\mathcal{\tilde{G}}(i\omega_{n})$ (right) with finite $N$.
+The top panels show absolute values of the PIMC data (solid) and the corresponding Drude value (dotted).
+The bottom panels shows the relative differences $\Delta\mathcal{\tilde{G}}(\tau)$ (solid lines) and $\Delta\mathcal{\tilde{G}}(i\omega_{n})$, including statistical uncertainties (dotted lines).
+However, let us first consider systematic merits of accuracy for the PIMC simulation based on energies.
 Finite-size effects of the energies are studied by varying the number of particles $N$ in the simulation, while keeping the density fixed.
 Results for kinetic and potential energies are presented in Fig.
-2 for up to $N=16$ and $\Theta=0.2$ at $r_{s}=4$.
+3 and for up to $N=16$ and $\Theta=0.2$ at $r_{s}=4$.
 Shown are also finite-size corrections [41, 36] ($\xi=0$), which rectify the energies toward the thermodynamic limit (TDL), according to:
 $K_{p}$ $=\frac{K_{N}}{N}+\frac{1}{N}\left(\frac{\omega_{p}}{4}+\frac{5.264}{4\pi r_{s}^{2}(2N)^{1/3}}\right)$ (47)
-$V_{p}$ $=\frac{V_{N}}{N}+\frac{\omega_{p}}{4N},$ (39)
-where the corrected energies $K_{p}$ and $V_{p}$ are further multiplied with $\mathrm{tanh(\omega_{p}\beta)}$.
+$U_{p}$ $=\frac{U_{N}}{N}+\frac{\omega_{p}}{4N},$ (48)
+where the corrected energies $K_{p}$ and $U_{p}$ are further multiplied with $\mathrm{tanh}(\omega_{p}\beta)$.
 The correlation energy is getting saturated with $N$, which also influences the kinetic energy.
 Simulation with $N>16$, or a reliable extrapolation to the TDL, would be needed to resolve the statistically significant finite-size bias of the energies.
-Also, clearly the Ewald summation with $k_{c}r_{c}>10$ is required to eliminate statistically significant bias from treating periodic images.
-Figure 3: Scaling of $\tilde{G}_{\mu\mu}(\tau)$ (left) and $\tilde{G}_{\mu\mu}(i\omega_{n})$ (right) with finite NN.
-The top panels show absolute values of the PIMC data (solid) and the corresponding Drude value (dotted).
-The bottom panels shows the relative differences $\Delta\tilde{G}_{\mu\mu}(\tau)$ (solid lines) and $\Delta\tilde{G}_{\mu\mu}(i\omega_{n})$, including statistical uncertainties (dotted lines).
+Also, clearly the Ewald summation with $r_{c}k_{c}>10$ is required to eliminate statistically significant bias from treating periodic images.
+Details of the Ewald parameters and a table of the energy values is found in the SI.
 To make physical benchmarks, there is obvious systematic bias related to the quantum statistics, i.e., the lack of Fermion exchange in the simulation.
@@
 Remaining bias ensues from using a finite time-step $\Delta\tau$ to treat higher-body correlations (beyond pair approximation) and periodic images.
-As the higher-body correlations intensify with density, we assume for this work a choice of the time-step according to $\Delta\tau=ars^{3}$, where $a=0.001$ Ha-1.
-The finite-size effects of the dipole polarizability are more topical: The scaling of $\tilde{G}_{\mu\mu}(\tau)$ and $\tilde{G}_{\mu\mu}(i\omega_{n})$ versus $N$ is plotted in Fig.
-3 for $k_{c}r_{c}=15.0$.
-Overall, the effects of $\tilde{G}_{\mu\mu}$ from PIMC versus Drude due to finite size are negligible regardless of $N$.
+As the higher-body correlations intensify with density, we assume for this work a choice of the time-step according to $\Delta\tau=ar_{s}^{3}$, where $a=0.001$ Ha-1.
+All these effects are elaborated and further discussed in the SI.
+The finite-size effects of the dipole polarizability are more topical: The scaling of $\mathcal{\tilde{G}}(\tau)$ and $\mathcal{\tilde{G}}(i\omega_{n})$ versus $N$ is plotted in Fig.
+4 for $r_{c}k_{c}=15.0$ at $r_{s}=4.0$ and $\Theta=0.2$.
+However, the finite-size effects of $\mathcal{\tilde{G}}$ due to $N$ between PIMC versus Drude are apparently negligible at this density.
 Nevertheless, we will use $N=16$ for the remainder of this work.
-The effects due to finite time-step and the Ewald cutoff parameter $k_{c}r_{c}$ are also statistically insignificant.
-The lack of finite-size effects in $\tilde{G}_{\mu\mu}$ is in stark contrast with the energies, which are affected by the correlation and, although absent here, exchange.
-The indifference to the number of particles is therefore a manifestation of a greater principle: the trajectories of an ensemble of free particles in a periodic setting sample the phase space in collective harmony.
-The dipole autocorrelation estimator only measures the collective state and distributes it evenly for each particle.
-Another way of seeing this is that the low-momentum states of the plasma are occupied and cannot be used to excite the dielectric response of individual particles.
-Curiously, the cases of interacting (PIMC) and non-interacting (Drude) come out effectively the same.
-Figure 4: On the left, $\tilde{G}_{\mu\mu}(\tau)$ and $\Delta\tilde{G}_{\mu\mu}(\tau)$ with variable $\Theta$ at $r_{s}=4$.
-On the right, $\tilde{G}_{\mu\mu}(i\omega_{n})$ and $\Delta\tilde{G}_{\mu\mu}(i\omega_{n})$ with variable $\Theta$ at $r_{s}=4$.
+The effects due to finite time-step and the Ewald cutoff parameter $r_{c}k_{c}$ are also statistically insignificant and only further discussed in the SI.
+Figure 5: On the left, $\mathcal{\tilde{G}}(\tau)$ and $\Delta\mathcal{\tilde{G}}(\tau)$ with variable $\Theta$ at $r_{s}=4$.
+On the right, $\mathcal{\tilde{G}}(i\omega_{n})$ and $\Delta\mathcal{\tilde{G}}(i\omega_{n})$ with variable $\Theta$ at $r_{s}=4$.
 Solid lines represent the absolute data and dotted lines the statistical uncertainties.
-Finally, we study how $\tilde{G}_{\mu\mu}(\tau)$ depends on physical parameters, the finite temperature and density.
+Figure 6: On the left, a comparison of $\mathcal{\tilde{G}}(\tau)$ versus $\tau$ from PIMC (solid black line) and the Drude model with variable values for $\Gamma$.
+On the right, the rescaled goodness measure $\eta$ is plotted against $\Gamma$ based on variables temperatures and $r_{s}=4$.
+Dotted lines are linear fits to the data.
+The absence of finite-size effects in $\mathcal{\tilde{G}}$ is attributed to the lack of many-body effects, overall, in the chosen parameter ranges.
+This is in stark contrast with the energies, indicating that the dipole polarizability is indeed much less sensitive to the dynamic correlation.
 In Fig.
-4 we plot $\tilde{G}_{\mu\mu}(\tau)$ and $\tilde{G}_{\mu\mu}(i\omega_{n})$ versus $\Theta$ with $r_{s}=4$ and $N=16$.
-The plotting axes of $\tilde{G}_{\mu\mu}(\tau)$ are rescaled to $\tau/\beta$, emphasizing that the absolute values of $\tilde{G}_{\mu\mu}(\tau=\beta/2)$ decrease with the inverse temperature.
-Consequently, the Matsubara integral $\tilde{G}(i\omega_{n}=0)$ goes down with $\beta$, but otherwise the Matsubara series $\tilde{G}(i\omega_{n}>0)$ follow the same curve, only with different spacings.
-Clearly, the PIMC results match the Drude model within statistical uncertainty of $2\%$ or less.
+5 we plot $\mathcal{\tilde{G}}(\tau)$ and $\mathcal{\tilde{G}}(i\omega_{n})$ versus $\Theta$ with $r_{s}=4$ and $N=16$.
+The plotting axes of $\mathcal{\tilde{G}}(\tau)$ are rescaled to $\tau/\beta$, emphasizing that the absolute values of $\mathcal{\tilde{G}}(\tau=\beta/2)$ decrease with the inverse temperature.
+Consequently, the Matsubara integral $\mathcal{\tilde{G}}(i\omega_{n}=0)$ goes down with $\beta$, but otherwise the Matsubara series $\mathcal{\tilde{G}}(i\omega_{n}>0)$ follow the same curve, only with different spacings.
+We present values of $\eta$ from Eq.
+(46) in Table 1 for various densities and temperatures.
+The PIMC results match the Drude model in all cases within at least $1\%$ accuracy.
+On the positive side, this means that the PIMC method is robust and accurate, albeit redundant, on these parameter ranges relevant for simulating metal structures.
+Again, the Drude response is also valid and unshaken by the account of Coulomb interactions, despite their effects on the energies.
+Yet, toward higher $r_{s}$ and lower $\Theta$ the value of $\eta$ is statistically different from 0, if only modestly, indicating that $\mathcal{\tilde{G}}(\tau)$ from PIMC is so much larger that $\mathcal{\tilde{G}^{D}}(\tau)$.
+Effects of the dynamic correlation are expected to intensify toward Wigner crystallization, only expected at much lower densities with $r_{s}\sim 10^{2}$, and lower temperatures [42].
 In Figure 6 we discuss how the match depends on the Drude relaxation parameter $\Gamma$.
-When $\Gamma$ is increased, the Drude correlation function $\tilde{G}_{\mu\mu}(\tau)$ loses magnitude.
+When $\Gamma$ is increased, the Drude correlation function $\mathcal{\tilde{G}}(\tau)$ loses magnitude.
 The phenomenological explanation is that the scattering rate of the charge carriers from ions increases with $\Gamma$, and thus, the spatial diffusion slows down.
@@
 This is presented in the lower panel of Fig.
-5, where linear fits approach $\eta\rightarrow 0$ when $\Gamma\rightarrow 0$.
+6, where linear fits approach $\eta\rightarrow 0$ when $\Gamma\rightarrow 0$.
 The relative error $\eta$ scales linearly with $\Gamma$ and $\beta$, and so, curves representing $\eta/\beta$ are overlaid within statistical uncertainties.
+For example, observed scattering constants for copper, silver and gold are, respectively, $3.5\times 10^{-3}$, $7.8\times 10^{-4}$ and $2.6\times 10^{-3}$ [4] in atomic units.
 With this consideration we want to suggest a future possibility of estimating $\Gamma$ from PIMC simulations, provided a reasonable but yet unknown sampling method of the scattering process.
 As seen from all the presented results, such a simulation could be easily validated, while providing first-principles insights to the Drude scattering.
-Furthermore, using a finite-$\Gamma$ simulation such as that could also lend itself for studying nonlinear optical phenomena of Drude-like materials on a quantum mechanical level.
-Figure 5: On the left, a comparison of $\tilde{G}_{\mu\mu}(\tau)$ versus $\tau$ from PIMC (solid black line) and the Drude model with variable values for $\Gamma$.
-On the right, the rescaled goodness measure $\eta/\beta$ is plotted against $\Gamma$ based on variables temperatures and $r_{s}=4$.
-Dotted lines are linear fits to the data.
+Furthermore, using a finite-$\Gamma$ simulation such as that could also lend itself for studying non-linear optical phenomena of Drude-like materials on a quantum mechanical level.
+On the other hand, since finite $\Gamma$ and correlation effects anticipated at high $r_{s}$ both increase $\eta$, they cannot compensate each other.
+Therefore, it remains to be established, how the many-body effects should be incorporated into an effective and predictive first-principles Drude model.
+$r_{s}$
+$\Theta$ 2.0 4.0 8.0 12.0
+0.1 0.2(0.7) 0.2(0.5) 0.9(0.5) 0.6(0.7)
+0.2 -0.0(0.5) -0.1(0.3) 0.4(0.2) 0.4(0.3)
+0.4 0.0(0.2) -0.1(0.2) 0.13(11) 0.3(0.2)
+0.8 0.1(0.2) -0.1(0.2) 0.12(11) 0.0(0.2)
+Table 1: The match parameter $\eta$ with variable $r_{s}$ and $\Theta$ using $N=16$, $r_{c}k_{c}=15$ and $\Gamma=0.0001$.
 V Summary and outlook
 We have investigated a PIMC estimator for the dipole polarizability of periodic quantum gases and plasmas in the long-wavelength limit, i.e., the optical regime.
-We have bench marked the estimator with quantum Coulomb plasma in a range of finite densities ($r_{s}=2\ldots 8$ a0) and finite temperatures ($\Theta=0.1\ldots 0.8$ and up to $T=10^{6}$ Kelvin).
-The PIMC results are compared to the Drude model in the imaginary domain, where they match within statistical limits and $\Delta\tilde{G}_{\mu\mu}<2\%$, when the Drude damping coefficient $\Gamma$ is small.
-That is, the PIMC results for $\tilde{G}_{\mu\mu}$ show no statistically significant effects due (reasonable) simulation parameters, such as finite time-step, periodic images or finite number of particles, nor physical parameters such as finite density and temperature.
-This corroborates that the Drude spectrum is suitable for describing the linear optical response (i.e.
-the long-wavelength limit) regardless of the coupling strength.
-Perhaps some minor methodological, systematic biases would have surfaced with generous statistical sampling, but this does not appear as a worthy goal.
-In summary: due to the collective motion of many interacting particles in a flat potential, all results come out consistent, effectively, with a free particle.
-This physical conclusion may not be overly interesting, but it firmly establishes that the numerical method, including the complete post-processing and analysis framework, is valid and robust.
-To this end, we have demonstrated special sampling approaches (see Sec.
-II.4), to allow accurate simulation even with a low particle count.
+We have bench marked the estimator with quantum Coulomb plasma in a range of finite densities ($r_{s}=2\ldots 12$ a0) and finite temperatures ($\Theta=0.1\ldots 0.8$ and up to $T=10^{6}$ Kelvin).
+The PIMC results are compared to the Drude model in the imaginary domain, where they match within statistical limits and $\Delta\mathcal{\tilde{G}}<1\%$, when the Drude $\Gamma$ parameter is small.
+That is, the PIMC results for $\mathcal{\tilde{G}}$ show no statistically significant effects due (reasonable) simulation parameters, such as finite time-step, periodic images or finite number of particles.
+Perhaps some minor methodological biases would have surfaced with generous statistical sampling, but this does not appear as a worthy goal.
+The physical parameters, finite density and temperature, show almost no effects due to many-body correlation in the chosen parameter ranges.
+Hints of correlation effects make exceptions at lower densities $r_{s}\geq 8$, i.e., higher coupling constants.
+On the one hand, this corroborates that the Drude spectrum is effective in describing free charge carriers at room temperatures and metallic densities.
+On the other hand, revisiting higher coupling constants with proper quantum statistics makes for an intriguing benchmarking prospect.
+Even in the absence of many-body effects, it is somehow fascinating to see reduction of the collection of interacting particles into an effectively free particle and a historical, phenomenological model.
+While this is theoretically anticipated, the numerical side conceals many hurdles which are now firmly established.
+To this end, we have demonstrated special post-processing, analyses and sampling approaches (see Sec.
+II.4) to allow accurate simulation even with a low particle count.
 These practices form a baseline for numerous further developments:
+$T$ (K) $\Gamma$ $\eta$ (%)
+1000.0 0.0001 1.1(1.1)
+0.001 4.3(1.2)
+0.01 35(2)
+0.1 264(5)
+10000.0 0.0001 -0.1(0.3)
+0.001 0.2(0.3)
+0.01 3.5(0.3)
+0.1 33.9(0.3)
+100000.0 0.0001 0.06(10)
+0.001 0.04(10)
+0.01 0.36(10)
+0.1 3.62(11)
+1000000.0 0.0001 0.07(14)
+0.001 0.02(14)
+0.01 0.05(14)
+0.1 0.39(14)
+Table 2: Model matching parameter $\eta$ with statistical uncertainties versus Drude $\Gamma$ parameter at variable fixed temperatures.
 i) The Fermion spectrum.
-The treatment of signed permutations should be considered in the future, to unfold the true electron gas behaviour, including complexities in the statistical performance, sampling acceleration and path winding.
-ii) Confined spectrum.
+The physically accurate treatment of signed permutations should be considered in the future.
+While this presents challenges with the statistical performance, sampling acceleration and path winding, it also provides an accurate benchmark of quantum many-body corrections in low temperatures and densities.
+ii) Inversion to real domain.
+Where applicable, migration of numerical many-body correction from the integral transformations such as Eqs.
+(18) or (20) is a formidable challenge.
+The maximum entropy approaches [27, 25, 21] have been developed, but they work better for sharp Lorentzian peaks than the broad Drude-like spectrum [24, 43].
+Finding analytical models and special approaches remains an open challenge.
+iii) Confined spectrum.
 Treating interacting particles in environments less regular than a flat potential (e.g.
-weak confinements, optical cavities or ionic plasma) should introduce non-trivial thermal and quantum effects for example with respect to the Drude–Lorentz model.
-iii) Inverse Laplace transform.
-Special methods to invert Eqs.
-(16) or (17) could be studied to migrate possible quantum corrections into effective models.
-The MaxEnt method is not the best suited for the broad, Drude-like spectra [23, 41] of this work, but in the presence of resonance peaks, it could be.
-iv) Higher-order response properties.
+weak confinements, optical cavities or ionic plasma) should introduce non-trivial thermal and quantum effects for example with respect to the Drude–Lorentz model, characterized with resonance frequencies.
+iv) Adsorption on surfaces.
+Dispersion interactions (e.g.
+van der Waals and Casimir forces) between atoms and surfaces can be expressed with the instantaneous multipole fluctuations, i.t., polarizabilities in the Matsubara frequency domain [44].
+As the PIMC method treats them fluently [29, 30], we may now also use it to inform effective adsorption models for Drude-like matter, such as HEG [45].
+v) Higher-order response properties.
 The response functions beyond the dipole order could be more sensitive to thermal quantum many-body effects.
 Their static and dynamic variants, such as $\chi^{(3)}$ of the Drude-like materials, could inform effective models [5] and applications, such as ENZ materials [46].
+VI Supplemental information
+The Supplemental information contains further benchmark figures of accuracy.
 Acknowledgements.
@@
 JT acknowledges Ilkka Kylänpää for his support in the implementation and assessment of periodic potentials.
+Appendix A Reduction of Lindhard theory to Drude model in the long-wavelength limit
+The Lindhard function in 3 dimensions, first mentioned in Sec. I, is given by [11]:
+$\chi_{0}(\textbf{q},\omega)=\frac{1}{L^{3}}\sum\frac{f_{\textbf{k}+\textbf{q}}{-f_{\textbf{k}}}}{E_{\textbf{k}+\textbf{q}}-E_{\textbf{k}}-\hbar(\omega+i\delta)},$ (49)
+where $\chi_{0}(\textbf{q},\omega)$ is the linear response to an incoming electric field with momentum q and frequency $\omega$, k and $E=\hbar^{2}k^{2}/2m_{e}$ are, respectively, the allowed momenta and associated energies in the electron gas, $m_{e}$ is the electron mass, $f_{\mathbf{k}}$ is the Fermi-Dirac distribution, and $\delta$ is an infinitesimal term associated to adiabatic switching-on.
+In the case of $G=0$, the response given by Eq.
+(1) is given by
+$\begin{split}\chi(\textbf{q},\omega)=&-v_{q}\chi_{0}(\textbf{q},\omega)\\ =&-\frac{v_{q}}{L^{3}}\sum\frac{f_{\textbf{k}+\textbf{q}}{-f_{\textbf{k}}}}{E_{\textbf{k}+\textbf{q}}-E_{\textbf{k}}-\hbar(\omega+i\delta)},\end{split}$ (50)
+where $v_{q}$ is the reciprocal Coulomb potential.
+It is well known that this susceptibility approaches the Drude model with $\Gamma=0$ by taking the limits of $q\rightarrow 0$ (known as the long-wavelength limit[6, 7]) and $\delta\rightarrow 0$ [47].
+The following procedure recovers the Drude model with finite $\Gamma$, as define by Eq.
+(36), by considering a finite $\delta$.
+The numerator of the sum in Eq. 50 can be approximated as
+$f_{\textbf{k}+\textbf{q}}-f_{\textbf{k}}\simeq\textbf{q}\cdot\nabla_{\textbf{k}}f_{\textbf{k}},$ (51)
+while the denominator can be written as
+$\begin{split}E_{\textbf{k}+\textbf{q}}-E_{\textbf{k}}&=\frac{\hbar^{2}}{2m_{e}}(k^{2}+2\textbf{k}\cdot\textbf{q}+q^{2})-\frac{\hbar^{2}k^{2}}{2m_{e}}-\hbar(\omega+i\delta)\\ &\simeq\frac{\hbar^{2}\textbf{k}\cdot\textbf{q}}{m_{e}}-\hbar(\omega+i\delta).\end{split}$ (52)
+Substituting these approximations in Eq. 50 leads to:
+$\chi(\textbf{q},\omega)\simeq-\frac{v_{q}}{L^{3}}\sum_{\textbf{k}}\frac{\textbf{q}\cdot\nabla_{\textbf{k}}f_{\textbf{k}}}{\frac{\hbar^{2}\textbf{k}\cdot\textbf{q}}{m_{e}}-\hbar(\omega+i\delta).}$ (53)
+Taking the factor $\hbar(\omega+i\delta)$ outside of the sum and approximating the denominator with its Taylor expansion results in
+[…]
+In Eq.
+54, the term proportional to $\sum_{\textbf{k}}\nabla_{\textbf{k}}f_{\textbf{k}}$ vanishes and the remaining term can be solved by partial integration [47].
+Then, the sum $\sum_{\textbf{k}}f_{\textbf{k}}$ corresponds to the number of electrons $N$.
+Finally, the fact that $v_{q}=e^{2}/|\textbf{q}|^{2}\varepsilon_{0}$ is used in the last step.
+At this point, the dependency on $q$ is dropped and the quantity $\omega^{2}_{p}=(ne^{2}/\varepsilon_{0}m_{e})$ is recognized as the plasma frequency with number density $n=N/L^{3}$.
+Using these definitions leads to
+$\chi(\omega)=-\frac{\omega_{p}^{2}}{(\omega+i\delta)^{2}}.$ (55)
+Finally, by defining $\delta=\Gamma/2$ and considering $\delta^{2}\approx 0$ , Eq. 37 is recovered
+$\chi^{D}(\omega)=-\frac{\omega_{p}^{2}}{\omega^{2}+i\omega\Gamma}.$ (56)
+We point that the fact that both $\delta$ in the Lindhard function and $\Gamma$ in the Drude model are infinitesimally small is the reason why we observe the best match in our results for small $\Gamma$, as seen in Table 2.